### v1

# TS-PEFT: Token-Selective Parameter-Efficient Fine-Tuning with Learnable Threshold Gating

Abstract
In the field of large models (LMs) for natural language processing (NLP) and computer vision (CV), Parameter-Efficient Fine-Tuning (PEFT) has emerged as a resource-efficient method that modifies a limited number of parameters while keeping the pretrained weights fixed. This paper investigates the traditional PEFT approach, which applies modifications to all position indices, and questions its necessity. We introduce a new paradigm called Token-Selective PEFT (TS-PEFT), in which a function S selectively applies PEFT modifications to a subset of position indices, potentially enhancing performance on downstream tasks. Our experimental results reveal that the indiscriminate application of PEFT to all indices is not only superfluous, but may also be counterproductive. This study offers a fresh perspective on PEFT, advocating for a more targeted approach to modifications and providing a framework for future research to optimize the fine-tuning process for large models.
1 Introduction
Large models (LMs) have achieved significant progress in natural language processing (NLP) yao2024survey, computer vision (CV) croitoru2023diffusion; liu2024sora, and many other fields. However, due to their large parameter scales, directly fine-tuning all parameters is often impractical in terms of computational and storage overhead. To address this issue, parameter-efficient fine-tuning (PEFT) has gradually become a mainstream paradigm for adapting pretrained large models: by freezing the base model network and introducing only a small number of trainable parameters, downstream customization can be accomplished efficiently hu2023llm; hanparameter.
Various PEFT methods have demonstrated strong performance across a wide range of downstream tasks. Among them, prompt-based approaches (such as Prefix-Tuning li2021prefix and Prompt-Tuning lester2021power) adjust model behavior by learning additional input tokens, while low-rank adaptation (LoRA) hu2022lora and its variants (e.g., AdaLoRA zhang2023adaptive, DoRA liu2024dora, VeRA kopiczkovera) update internal representations by injecting a small number of trainable parameters or low-rank components into the base model. These approaches significantly reduce the number of trainable parameters while typically maintaining performance comparable to, or even approaching, full-parameter fine-tuning.
Most existing PEFT methods, however, implicitly rely on a key assumption: once a layer is selected as the target module for PEFT, all token positions passing through this module should receive a trainable modification. Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is typically written as:
$h_{i}=W_{0}x_{i}+\operatorname{M}(x_{i}),$ (1)
where $\operatorname{M}(\cdot)$ is a PEFT module such as LoRA. This “update-all-positions” design ignores the fact that different token positions contribute unequally to the downstream objective.
Figure 1: Comparison between standard PEFT and our TS-PEFT framework.
In practice, we observe that, for already strong pretrained models, the deviation induced by PEFT is extremely small on many token positions, while only a small subset of positions exhibit substantial and task-relevant changes. In other words, most PEFT updates are effectively redundant at the token level. Uniformly applying modifications to all positions not only injects unnecessary noise and potential interference, but also increases computation, especially when multiple PEFT modules are stacked. This motivates a natural question: can PEFT be made selective over token positions instead of blindly updating every one of them?
Our empirical findings suggest a negative answer. We propose a simple yet effective framework, Token-Selective PEFT (TS-PEFT), which enables token-level selective parameter-efficient fine-tuning. Instead of always applying $\operatorname{M}(x_{i})$ to every position, TS-PEFT introduces a binary gating function $S$ at each token position that decides whether to add the PEFT update or to keep the base model output unchanged. Intuitively, positions whose PEFT updates are small relative to the frozen output remain untouched, while positions with larger, more impactful updates retain their PEFT modifications, yielding a token-level sparse pattern of PEFT updates.
In the experimental section, we conduct systematic evaluations on representative models including LLaMA-3-8B, LLaVA-1.5-7B, and DeBERTaV3-base hedebertav3, covering commonsense reasoning, visual instruction tuning, and natural language understanding. The results show that, on top of strong PEFT baselines such as LoRA, DoRA, AdaLoRA, and Adapter, TS-PEFT consistently matches or surpasses their performance while applying updates to only about 40–60% of token positions. Moreover, the learned token-level sparsity patterns can be aggregated into module-level importance indicators, which enable selecting fewer yet more critical tuning modules that even outperform full-module fine-tuning under comparable parameter budgets. In summary, our contributions are as follows:
  • We point out and examine a neglected aspect of standard PEFT: once a layer is chosen as a target module, all token positions are updated uniformly, even though only a small fraction of positions contribute meaningfully to the downstream objective. Our empirical analysis of token-level sparsity shows that many PEFT updates induce negligible deviation from the frozen backbone, revealing substantial redundancy along the sequence dimension and motivating selective updates.
  • We propose TS-PEFT, which introduces a token-level binary gating function on top of existing PEFT modules. By formulating the learning of layer-wise thresholds as a sparsity-regularized proximal optimization problem, deriving a practical approximate gradient, and applying Adam-style moment estimation, TS-PEFT enables stable end-to-end training of token-level gates jointly with PEFT parameters. The framework is agnostic to the specific low-rank parameterization and can be used as a plug-in module for a variety of PEFT methods.
  • On three representative benchmarks, TS-PEFT consistently matches or surpasses strong PEFT baselines, while updating only 40%-60% token positions. Moreover, the learned token-level sparsity patterns can be aggregated into module-wise importance scores, enabling selective tuning of fewer but more critical modules that can outperform full-module fine-tuning under comparable parameter budgets.
2 Related Work
Low-Rank Adaptation. Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while preserving downstream performance. LoRA hu2022lora introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models; after training, these matrices can be merged back into the base model without incurring additional inference cost. Building on this idea, AdaLoRA zhang2023adaptive adaptively redistributes the effective rank across different layers to better meet task-specific requirements, DoRA liu2024dora refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA kopiczkovera leverages shared low-rank bases to further reduce per-task parameter overhead. However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update—that is, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension. This may introduce unnecessary redundant modifications and fails to explicitly characterize or exploit potential token-level redundancy.
Sparsity. Sparsity-guided methods have been widely used to improve the efficiency and interpretability of large language models. By reducing the number of active parameters and concentrating computation on core components, these approaches not only make models more efficient but also offer clearer insights into their decision-making processes. Mixture-of-experts (MoE) models, exemplified by Switch Transformers fedus2022switch and Mixtral 8x7B jiang2024mixtral, employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters, and several recent works have combined MoE architectures with LoRA-style adapters to further enhance performance chen2024llava; liu2024moe; zadouripushing . Another typical use of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets ma2023llm; xubesa; xiasheared. tan2024sparsity further introduce sparsity-guided techniques as a tool for interpreting the behavior of LLMs. However, most existing studies focus on parameter-wise sparsity, i.e., deciding which weights or modules to activate or prune. In contrast, our work explores token-level sparsity in the context of PEFT: instead of sparsifying parameters, we selectively decide which token positions should receive PEFT updates, revealing and leveraging redundancy along the sequence dimension.
3 Method
This section presents the overall theoretical framework and optimization process of TS-PEFT. We first formalize the LoRA-style PEFT procedure and introduce a token-level control function $\operatorname{S}$ to determine whether a PEFT update should be applied at each position. Then, we formulate the learning of the scalar threshold $\tau$ in $\operatorname{S}$ as an optimization problem with a sparsity regularization term. Next, to address the non-differentiability of $\operatorname{S}$ with respect to $\tau$ and the issue that gradients are nearly zero at most positions, we derive a computable approximate gradient. Finally, we combine this approximate gradient with momentum and adaptive learning-rate mechanisms so that it can be seamlessly integrated into the Adam optimizer, and we provide the overall algorithmic description of the TS-PEFT training process.
3.1 Token-level Selective PEFT Formulation
As shown in Eq.(1), once a layer is selected as a PEFT target layer, all token positions passing through that layer will uniformly receive the increment introduced by $\operatorname{M}(x_{i})$. To break this assumption, we introduce a token-level gating function $\operatorname{S}$ on top of Eq.(1) to determine whether a PEFT update is actually applied at each position. Specifically, the output $h_{i}$ at position $i$ is rewritten as:
$h_{i}=\begin{cases}W_{0}x_{i}&\text{if }\operatorname{S}(W_{0}x_{i},\operatorname{M}(x_{i}))=0,\\ W_{0}x_{i}+\operatorname{M}(x_{i})&\text{if }\operatorname{S}(W_{0}x_{i},\operatorname{M}(x_{i}))=1.\end{cases}$ (2)
Here, $\operatorname{S}(\cdot)\in\{0,1\}$ serves as a gating variable: when $\operatorname{S}=1$, the update is enabled at that position, and when $\operatorname{S}=0$, the position retains the base model output. To instantiate Eq.(2), we focus on the relative magnitude $r_{i}$ of the PEFT update with respect to the base model output at each position $i$: $r_{i}=\|\operatorname{M}(x_{i})\|_{2}/\|W_{0}x_{i}\|_{2}$. This is because the feature norm values at different layers vary significantly. The relative values enable each layer to adaptively adjust according to its own weight change ratio. Intuitively, a larger $r_{i}$ means that the PEFT update induces a stronger perturbation relative to the base model output at that position and is more likely to be task-relevant, whereas a small $r_{i}$ indicates that the modification is minor and can be treated as negligible. Therefore, we introduce a learnable scalar threshold $\tau$ for each target layer and define the control function accordingly:
$\operatorname{S}(W_{0},\operatorname{M}(x_{i}))=\begin{cases}0&\text{if }r_{i}<\tau,\\ 1&\text{if }r_{i}\geq\tau.\end{cases}$ (3)
In other words, when the relative magnitude at a given position is not lower than $\tau$, we consider the update at that position worth retaining; otherwise, the update for that position is skipped. This results in a token-level sparse update pattern along the sequence dimension within each selected PEFT module, with the threshold $\tau$ directly controlling the overall sparsity level.
However, this threshold-based step gating introduces two training challenges: first, $\operatorname{S}$ is a non-smooth step function with respect to $\tau$, so standard backpropagation provides almost no useful gradients; second, we want $\operatorname{S}$ to activate PEFT updates only on a subset of token positions while keeping the remaining ones identical to the backbone, which requires explicit control of the overall sparsity level during threshold learning. The next subsections tackle these two issues via a proximal objective and a computable gradient approximation.
3.2 Proximal Optimization for the Threshold
Due to the training difficulties mentioned above, directly using first-order optimization methods cannot effectively update $\tau$. Therefore, to ensure that the threshold can be stably updated during training, we formulate its learning process as a proximal optimization problem, allowing it to follow the changes in the original task loss while explicitly favoring a sparser update pattern. Specifically, our objective function is defined as follows:
$\min\mathcal{L}(\tau)=\ell(\tau)+\lambda\sum_{i}\mathbbm{1}(r_{i}\geq\tau),$ (4)
where $\ell(\tau)$ denotes the loss of the target task, and the second term penalizes the number of excessively activated tokens to encourage a sparser update, $\mathbbm{1}(\cdot)$ is the indicator function, and $\lambda$ serves as a hyperparameter that balances the sparsity level.
However, since the indicator function is non-differentiable, we adopt proximal optimization techniques and introduce a quadratic regularization term in the threshold update to limit its fluctuation parikh2014proximal. Given the threshold $\tau^{k}$ at the $k$-th iteration, the update $\tau^{k+1}$ can be written as:
$\underset{\tau}{\mathrm{argmin}}\left[\ell(\tau)+\frac{\partial\ell}{\partial\tau}(\tau-\tau^{k})+\frac{1}{2\alpha_{k}}\left|\tau-\tau^{k}\right|^{2}+\lambda P(\tau)\right],$ (5)
where $\alpha_{k}$ is the step size. For convenience in the subsequent derivation, we set $\sum_{i}\mathbbm{1}(r_{i}\geq\tau)$ to be $P(\tau)$. It can be seen that this subproblem consists of three components: a first-order approximation of the loss at $\tau^{k}$, a proximal term that constrains the magnitude of the update, and a sparsity regularization term that favors fewer activated positions.
To solve Eq.(5), we need to compute $\partial\ell/\partial\tau$ and $\partial P/\partial\tau$. Thus, we first expand the output in Eq.(2) along the hidden dimension. Let the hidden representation of the $i$-th token at dimension $j\in\{0,\dots,H-1\}$ be denoted as:
$h_{ij}=\left[W_{0}x_{i}\right]_{j}+\left[\operatorname{M}(x_{i})\ \right]_{j}*\mathbbm{1}\left(r_{i}\geq\tau\right),$ (6)
where $H$ is the hidden size, and $\cdot]_{j}$ denotes the $j$-th component of a vector. By backpropagation, we finally get:
$\frac{\partial\ell}{\partial\tau}=\sum_{i}\mu_{i}\frac{\partial}{\partial\tau}\mathbbm{1}\left(r_{i}\geq\tau\right),$ (7)
$\frac{\partial P}{\partial\tau}=\sum_{i}\lambda\frac{\partial}{\partial\tau}\mathbbm{1}\left(r_{i}\geq\tau\right),$ (8)
where $\mu_{i}$ aggregates the overall influence of the PEFT update at position $i$ on the loss:
$\mu_{i}=\sum_{j=0}^{H-1}\frac{\partial\ell}{\partial h_{ij}}\left[M(x_{i})\ \right]_{j}.$ (9)
3.3 Gradient Approximation for the Step Function
In the previous subsection, we obtained the derivative expressions of the loss function and the sparsity penalty term with respect to the threshold $\tau$. However, all these derivatives depend on the partial derivative of the step gating $\mathbbm{1}(r_{i}\geq\tau)$ with respect to $\tau$, whose numerical value is almost zero everywhere, making it impossible to directly perform gradient updates. Therefore, to enable stable learning of the threshold through backpropagation, this section constructs a computable gradient approximation (based on the perspective of proximal optimization) to replace the non-differentiable derivative of the indicator function.
Algorithm 1 Forward and Backward
1:Input: $\operatorname{M}$, $\alpha_{k}$, $s$, $\lambda$, $\beta_{1}$, $\beta_{2}$, $\tau^{k}$.
2:Output: $\tau^{k+1}$, updated parameters of $\operatorname{M}$ .
3:Forward Pass:
4:  Forward as in Eq.(2).
5:Backward Pass:
6:  Step 1: Perform the usual backward step to update parameters of $\operatorname{M}$.
7:  Step 2:
8:   Compute $\mu_{i}$ as in Eq.(9).
9:   Compute $g_{k}$ as in Eq.(15).
10:   Update $\tau^{k+1}$ as in Eq.(20).
A straightforward approximation is to treat the gradient of the step function as a constant s,s>0-s,s>0, that is:
$\frac{\partial}{\partial\tau}\mathbbm{1}\left(r_{i}\geq\tau\right)\approx-s.$ (10)
Substituting this approximation into Eq.(7)-(9) yields the “rough” gradients of the loss term and the sparsity term with respect to the threshold:
$\frac{\partial\ell}{\partial\tau}$ $\approx-s\sum_{i}\mu_{i},$ (11)
$\frac{\partial P}{\partial\tau}$ $\approx-s\sum_{i}\lambda.$ (12)
However, we observe that Eq.(12) will always push $\tau$ upward, while the direction of Eq.(11) depends on the sign of $\mu_{i}$, making Eq.(10) a rather coarse approximation. To avoid unstable oscillation, we further introduce a consistency mask so that only tokens whose PEFT updates “should be activated” contribute to the threshold update. Concretely, we only use positions where the current gating decision agrees with their importance estimate ($\mathbbm{1}(\mu_{i}\geq 0)=\mathbbm{1}(r_{i}\geq\tau)$). Therefore, we revise Eq.(11) and Eq.(12) as follows:
$\frac{\partial\ell}{\partial\tau}=-s$ $\sum_{i}\mathbbm{1}\left[\mathbbm{1}\left(\mu_{i}\geq 0\right)=\mathbbm{1}\left(r_{i}\geq\tau\right)\right]\mu_{i},$ (13)
$\frac{\partial P}{\partial\tau}=-s\sum_{i}\mathbbm{1}\left(r_{i}\geq\tau\right)\lambda.$ (14)
The constraints can be interpreted as follows: whenever $\tau$ is either large enough ($\mu_{i}{\geq}0$, $r_{i}{<}\tau$) or small enough ($\mu_{i}{\leq}0$, $r_{i}{\geq}\tau$) to render the update of $\tau$ unnecessary, Eq.(13) zeros $\partial\ell/\ \partial\tau$; similarly, when $\tau$ is large enough ($r_{i}{<}\tau$, $\lambda$ is a positive constant), Eq.(14) nullifies $\partial P/\ \partial\tau$.
Combining both parts, the approximate gradient for updating the threshold can be unified as:
$g_{k}=\sum_{i}\left\{\mathbbm{1}\left[\mathbbm{1}\left(\mu_{i}\geq 0\right)=\mathbbm{1}\left(r_{i}\geq\tau\right)\right]\mu_{i}+\mathbbm{1}\left(r_{i}\geq\tau\right)\lambda\right\}.$ (15)
Finally, given the learning rate $\alpha_{k}$, the threshold is updated as $\tau^{k+1}=\tau^{k}+\alpha_{k}s\cdot g_{k}$, where $\alpha_{k}s$ can be regarded as the learning rate.
3.4 Adam-style Threshold Update and Overall Training Procedure
In practice, the approximate gradient $g_{k}$ for the threshold $\tau$ is highly noisy and varies in scale across training steps: the contributions of tokens change significantly between batches, and the magnitude of $g_{k}$ can fluctuate during optimization. Directly applying plain gradient descent to $\tau$ therefore tends to cause oscillation and slow or unstable convergence. To mitigate this, we update each threshold with an Adam-style rule that combines momentum and adaptive learning rateskinga2015method.
At the $k$-th iteration, we use the approximate gradient $g_{k}$ obtained in the previous subsection to update the first and second moments:
$m_{k}$ $\leftarrow\beta_{1}m_{k-1}+(1-\beta_{1})g_{k},$ (16)
$v_{k}$ $\leftarrow\beta_{2}v_{k-1}+(1-\beta_{2})g_{k}^{2},$ (17)
$\hat{m}_{k}$ $\leftarrow m_{k}/(1-\beta_{1}^{k}),$ (18)
$\hat{v}_{k}$ $\leftarrow v_{k}/(1-\beta_{2}^{k}),$ (19)
where $\beta_{1}$ controls the contribution of past gradients and $\beta_{2}$ governs the decay rate of squared gradients, while $\hat{m}_{k}$ and $\hat{v}_{k}$ denote the bias-corrected first- and second-moment estimates used in the Adam update. Finally, the threshold is updated as:
[…]
where $s$ is the gradient scaling factor, and $\epsilon$ prevents division by zero. In practice, this Adam-style update yields stable convergence of the thresholds under noisy gradients while preserving the desired sparsity level.
Hyperparams PIQA BoolQ HellaSwag WinoGrande SIQA OBQA ARC-e ARC-c Avg
LoRA $s=4\mathrm{e}{-5}$ $\lambda=4.5\mathrm{e}{-5}$ 88.5 64.2 95.3 85.5 80.8 85.4 90.3 79.9 83.7
TS-LoRA 88.6 70.1 95.5 84.4 82.3 85.8 90.1 79.4 84.5
Sparsity(%) 58.8 55.6 57.1 54.0 54.9 56.8 59.2 60.0 57.1
DoRA $s=4\mathrm{e}{-5}$ $\lambda=1\mathrm{e}{-5}$ 87.5 75.0 95.5 85.5 80.1 85.8 90.8 79.5 85.0
TS-DoRA 88.8 75.2 95.5 87.1 80.2 86.6 91.2 80.1 85.6
Sparsity(%) 50.0 47.9 50.2 47.1 47.6 48.3 49.6 50.0 48.8
AdaLoRA $s=4\mathrm{e}{-5}$ $\lambda=1\mathrm{e}{-4}$ 88.8 74.2 95.5 85.2 79.6 85.4 91.2 79.5 84.9
TS-AdaLoRA 88.4 75.5 95.8 85.6 80.9 86.2 90.1 79.9 85.3
Sparsity(%) 74.3 76.0 65.0 70.8 68.9 72.8 75.5 75.9 67.7
Table 1: Evaluation scores of LLaMA3.1-8B on CSR benchmarks; the last row in each block shows the sparsity of TS-PEFT.
Combining the above components, the training process of TS-PEFT can be summarized as follows: in each training step, during the forward pass, we perform gating based on each token’s relative update magnitude $r_{i}$ and the threshold $\tau$ of the corresponding layer, applying updates only at positions that satisfy the condition while keeping the backbone output unchanged at all other positions. Then, standard backpropagation is used to update all PEFT parameters, while the backbone parameters remain frozen throughout. On this basis, we compute $\mu_{i}$ to construct the approximate gradient $g_{k}$ for the threshold, and independently update the threshold $\tau$ of each layer using Adam-style first- and second-moment estimates. The full training procedure is presented in Algorithm 1.
4 Experiments
4.1 Experimental Setup
We evaluate TS-PEFT across three representative scenarios: natural language understanding (NLU), commonsense reasoning (CSR), and visual instruction tuning (VIT). The experimental setup covers the datasets, baseline models, and hyperparameter configurations to ensure reproducibility and comparability of the results.
Datasets and Models.
We select several representative benchmarks in the current PEFT and small-model fine-tuning literature. For CSR, we evaluate on LLaMA-3.1-8B dubey2024llama using PIQA, BoolQ, HellaSwag, WinoGrande, SIQA, OBQA, and ARC-e/ARC-c, covering typical language reasoning scenarios such as physical commonsense, Boolean question answering, narrative completion, and scientific QA liu2024dora. For VIT, we evaluate LLaVA-1.5-7B liu2023visual; liu2024improved on ScienceQA, POPE, MMBench, GQA, VisWiz, VQAv2, and VQAT, which are generative language tasks with open-ended answers conditioned on images and textual instructions, rather than fixed-choice classification. For NLU, we build on DeBERTaV3-base hedebertav3 and use the GLUE benchmark datasets wang2018glue. The GLUE tasks follow the official evaluation protocol: CoLA uses the Matthews correlation coefficient, STS-B uses correlation coefficients, and all other tasks use accuracy.
Baselines.
In our experiments, we select several mainstream parameter-efficient fine-tuning methods as baselines, covering two major categories: low-rank update methods and adapter-based methods. Specifically, LoRA, DoRA, and AdaLoRA are used to examine whether TS-PEFT can consistently provide additional gains under different low-rank update strategies, while Parallel Adapter hu2023llm is included to evaluate the method’s scalability on adapter-style architectures. For each baseline, we keep its original implementation, hyperparameters, and rank configuration unchanged, and apply TS-PEFT on top of it. This ensures that the comparison is not affected by factors unrelated to the baseline itself and that the results accurately reflect the impact of token-level selective updates.
Training details.
In all experiments, we freeze the backbone model and update only the PEFT parameters and the threshold $\tau$. We use the AdamW optimizer, where the PEFT parameters and $\tau$ share the same base learning rate, while the effective learning rate of $\tau$ is further modulated by the scaling factor s. For text tasks, we follow standard instruction-tuning configurations; for vision–language tasks, we adopt the original training pipeline of the corresponding models. Unless otherwise specified, we set $\alpha_{k}=1$, $\beta_{1}=0.9$, and $\beta_{2}=0.98$. Each main experiment is repeated at least three times with different random seeds, and we report the mean scores across runs. Additional hyperparameter settings are summarized in Appendix A.4.
4.2 Experimental Results
Hyperparams SQA POPE MMBench GQA VisWiz VQAv2 VQAT Avg
LoRA $s=4\mathrm{e}{-5}$ $\lambda=2\mathrm{e}{-6}$ 68.1 87.3 64.8 63.1 46.8 79.1 57.3 66.6
TS-LoRA 68.3 87.5 63.6 63.3 50.4 78.3 57.5 67.0
Sparsity(%) 67.6 59.3 62.0 59.3 59.5 59.2 62.2 61.3
DoRA $s=4\mathrm{e}{-5}$ $\lambda=1\mathrm{e}{-6}$ 67.5 87.4 65.5 63.0 50.5 78.6 57.0 67.1
TS-DoRA 69.4 88.0 64.7 62.4 54.2 77.8 55.9 67.4
Sparsity(%) 70.6 62.6 65.2 62.4 62.9 63.8 65.2 64.7
AdaLoRA $s=4\mathrm{e}{-5}$ $\lambda=5\mathrm{e}{-7}$ 68.0 87.3 64.3 61.7 48.9 78.3 56.9 66.5
TS-AdaLoRA 67.8 87.5 63.4 62.3 52.6 78.5 55.2 66.8
Sparsity(%) 60.0 50.9 54.1 50.9 51.4 50.9 53.9 53.2
Table 2: Evaluation scores of LLaVA-1.5-7B on VIT benchmarks; the last row in each block shows the sparsity of TS-PEFT.
Method MNLI CoLA QNLI RTE MRPC QQP SST-2 STS-B Avg
LoRA 89.9 69.2 94.1 87.1 90.4 92.2 95.3 91.8 88.8
TS-LoRA 89.9 70.3 94.2 88.1 91.2 92.0 95.8 91.5 89.1
Sparsity(%) 71.6 64.1 70.6 42.8 55.4 57.1 67.6 33.9 57.9
$\lambda$ 1e-5 6e-6 2e-5 1e-5 1e-5 2e-6 2e-5 8e-5 -
AdaLoRA 90.4 71.1 94.6 88.6 91.3 91.9 95.9 91.9 89.5
TS-AdaLoRA 90.5 71.3 94.5 89.0 91.6 91.9 96.1 91.6 89.6
Sparsity(%) 55.2 53.1 68.1 62.7 71.7 54.7 56.9 68.3 61.3
$\lambda$ 9e-6 5e-6 4.5e-7 2e-6 2e-8 5e-7 2e-6 5e-6 -
Adapter 90.4 72.1 94.2 86.6 90.0 92.1 95.6 91.4 89.1
TS-Adapter 90.1 72.0 94.1 86.8 91.4 91.8 96.6 91.3 89.3
Sparsity(%) 56.4 59.4 65.5 62.7 55.1 58.0 56.2 67.7 60.1
$\lambda$ 5e-7 4.5e-7 4.5e-7 2e-6 2e-8 5e-7 2e-6 5e-6 -
Table 3: Evaluation scores of DeBERTaV3-base on GLUE benchmarks; the “Sparsity(%)” rows show the sparsity of TS-PEFT, and the $\lambda$” rows list task-wise thresholds ( $s$ = 1e-4).
Commonsense Reasoning. Table 1 reports the results of various PEFT methods and their TS-PEFT variants on the CSR tasks. For a fair comparison, we set the rank of both LoRA and DoRA to 32, and configure AdaLoRA with an initial rank of 64 and a target rank of 32. Under the same rank settings, we observe that TS-LoRA, TS-DoRA, and TS-AdaLoRA achieve performance comparable to or better than their respective baselines on most datasets. This indicates that selectively skipping updates on a subset of tokens does not weaken downstream performance; instead, it can help mitigate the interference caused by redundant updates.
The table also reports the sparsity rate of TS-PEFT (i.e., the proportion of positions in the input sequence that are not modified by the PEFT output, as defined in Eq.(21)). Across different baselines and tasks, this proportion remains roughly within the 40%–60% range, meaning that fewer than half of the tokens actually receive PEFT updates. Even under this degree of sparsity, TS-PEFT is able to match or outperform the original baselines in commonsense reasoning, suggesting that standard PEFT exhibits substantial redundancy along the token dimension, and that threshold-based token-level selective updating provides a more efficient use of PEFT capacity.
$\text{Sparsity}(W_{0},\operatorname{M},X)=1-\frac{1}{T}\sum_{i}{\operatorname{S}(W_{0}x_{i},\operatorname{M}(x_{i}))}.$ (21)
Visual Instruction Tuning. For visual instruction tuning, we follow the training setup of liu2024dora, configuring LoRA and DoRA with rank 128 and setting the initial and target ranks of AdaLoRA to 256 and 128. Table 2 shows the results. Across all benchmarks, TS-LoRA, TS-DoRA, and TS-AdaLoRA yield consistent improvements of about 0.3–0.4 points, while operating under 50–65% token sparsity. This demonstrates that token-level redundancy is also present in multimodal generative tasks, and selective token updates remain effective, consistent with our observations in CSR.
Natural Language Understanding. The results are shown in Table 3, where the target ranks of LoRA and AdaLoRA are both set to 4. Overall, TS-LoRA and TS-AdaLoRA achieve average performance that is comparable to or slightly better than their respective baselines, consistent with the trends observed in the previous two task categories. To test whether our approach is tied to a specific low-rank parameterization, we further include Parallel Adapter as a non–LoRA-style PEFT baseline and apply TS-PEFT on top of it to obtain TS-Adapter, using a hidden size of 32. The results show that TS-Adapter performs no worse than the original Adapter, suggesting that TS-PEFT is not bound to LoRA-style weight structures, but can function as a general token-level gating layer that can be stacked on diverse PEFT modules.
TS-PEFT introduces token-level sparsity as a logical mask. Therefore, the current implementation does not provide wall-clock speedups. However, the theoretical computation reduction is significant. With specialized kernels support for sparse matrix additions in the future, TS-PEFT has the potential to translate this logical sparsity into physical latency reduction.
4.3 In-depth Analysis
Figure 2: Performance of LLaMA3.1-8B on CSR benchmarks as the percentage of selected modules varies for each PEFT method.
PIQA BoolQ HellaSwag WinoGrande SIQA OBQA ARC-e ARC-c Avg
S_low_50% 88.6 75.7 95.4 86.9 80.6 87.6 90.8 79.7 85.7
S_high_50% 88.2 62.2 92.2 85.2 80.2 86.2 90.4 79.9 83.0
norm_relative_50% 88.4 62.2 95.4 85.0 80.5 83.6 90.1 80.2 83.2
norm_abs_50% 88.4 52.7 85.8 83.9 78.7 86.2 89.6 79.0 80.5
random_50% 88.3 73.4 91.2 85.6 80.6 85.7 89.6 78.9 84.2
half_rank_50% 88.1 75.4 95.2 86.3 81.2 85.2 89.6 78.7 84.9
Table 4: The performance of LLaMA3.1-8B on the commonsense reasoning task using various selection methods.
Necessity of Adaptive Moment Estimation. To verify the effect of the adaptive momentum estimation module, we conduct ablation experiments by removing this component. In this case, the threshold $\tau$ exhibits noticeable oscillation and gradually diverges, and the training loss also fails to converge. On representative CSR tasks such as BoolQ, this instability leads to a clear degradation in dev performance compared with the full TS-PEFT variant (see Appendix A.2 for detailed curves and numerical results). These observations indicate that adaptive momentum estimation is essential for stabilizing the threshold update process in practice.
Module-wise Sparsity Distribution. In LLaMA3.1-8B, there are 32 layers in total. We compute the sparsity of q_proj, k_proj, and v_proj in each layer and take their average to obtain the sparsity distribution of TS-PEFT across modules (detailed results are provided in Appendix A.3). Overall, the sparsity of TS-LoRA and TS-AdaLoRA is relatively evenly distributed across layers, with standard deviations of 9.8 and 11.4, respectively. In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, showing a more “polarized” pattern. It should be noted that LoRA and AdaLoRA share the same weight structure despite having different rank allocation strategies, whereas DoRA adopts a distinct parameterization by decomposing weights into magnitude and direction. We hypothesize that this structural difference leads to the stronger selectivity observed in TS-DoRA at the module level, naturally raising the question: are the modules with lower sparsity more critical for downstream tasks?
Token-level Sparsity as a Module Importance Indicator. To examine whether token-level sparsity can be used to identify important modules, we design a set of selection experiments based on the sparsity patterns learned by TS-DoRA. After completing TS-DoRA fine-tuning, we rank all modules by their average token sparsity and construct two main settings under the same parameter budget: (1) S_low_50%, which fine-tunes only the 50% of modules with the lowest sparsity, and (2) S_high_50%, which fine-tunes only the 50% of modules with the highest sparsity.
We compare these sparsity-based selections against several alternatives: norm_relative_50% (selecting the 50% of modules with the largest average $\lVert\operatorname{M}(x_{i})\rVert_{2}/\lVert W_{0}x_{i}\rVert_{2}$), norm_abs_50% (the 50% with the largest average $\lVert\operatorname{M}(x_{i})\rVert_{2}$), random_50% (randomly selecting 50% of modules, averaged over three runs), and half_rank_50% (using all modules but halving the rank so that the total number of trainable parameters matches the 50% settings). The results in Table 4 show that S_low_50% not only significantly outperforms S_high_50% under the same parameter budget, but even surpasses fine-tuning all DoRA modules (whose mean score is 85.0). In contrast, norm-based selections, random choices, and half-rank full-module tuning all yield clearly inferior performance. This indicates that, in our setting, the token-level sparsity patterns learned by TS-DoRA provide a more reliable signal of module importance than simple norm statistics or random selection.
Effect of the Number of Target Modules. After using sparsity as the criterion for module importance ranking, we further examine the effect of varying the proportion of target modules while keeping the ranking fixed: we select different percentages of modules in ascending order of sparsity and fine-tune them. The results are shown in Figure 2. For AdaLoRA, performance remains largely unaffected as the proportion of target modules increases. However, for the others, “more is better” does not hold: DoRA achieves its best performance when using only 50% of the modules, LoRA reaches its optimal performance at 80%, and even using only 20% of the modules yields results close to the optimum.
In addition, LoRA and DoRA exhibit clear performance cliffs at certain points: for DoRA, performance drops by about 6.6% when increasing the target module proportion from 50% to 60%; for LoRA, increasing the proportion from 90% to 100% results in a drop of about 1.3%. This phenomenon indicates that jointly training modules with higher sparsity may negatively interfere with overall performance. In contrast, AdaLoRA is less sensitive to such “low-importance modules,” as its dynamic rank allocation redistributes capacity across modules.
5 Conclusion
This paper revisits parameter-efficient fine-tuning from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers. We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 40–60% of token positions, yet consistently matches or surpasses strong baselines. Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a basis for future work on more efficient sparse fine-tuning and inference.
Appendix A Appendix
A.1 LLM Usage Statement
During the preparation of this manuscript, the authors utilized a large language model (LLM) exclusively for grammar and language polishing. All technical content, scientific claims, and experimental results were conceived, derived, and verified solely by the authors.
A.2 The training loss progression with and without the Adaptive Moment Estimation
Figure 3 illustrates the training dynamics of a single TS-PEFT layer (the v_proj in layers.20.self_attn of LLaMA3.1-8B) with and without Adam-style momentum on the threshold $\tau$. In both subfigures, the blue curve shows the training loss, while the red dashed curve tracks the corresponding threshold value over training steps (right-hand axis). When momentum is enabled (Figure 3(a)), the threshold increases smoothly and gradually saturates around a stable value, while the loss decreases rapidly at the beginning and then oscillates mildly around a low level. This indicates that $\tau$ can adapt to the task in a stable manner without causing large fluctuations in the optimization trajectory.
Figure 3: Training loss progression for layer base_model.layers.20.self_attn.v_proj of LLaMA3.1-8B.
In contrast, when we remove the adaptive moment estimation and update $\tau$ using plain gradient descent (Figure 3(b)), the threshold trajectory becomes highly oscillatory and eventually drifts to a much larger range (above 0.6). The loss curve also exhibits large, persistent oscillations and fails to converge to a similarly low level as in Figure 3(a). This unstable behavior at the layer level is consistent with the degradation observed on downstream CSR benchmarks such as BoolQ: the dev accuracy drops noticeably compared with the full TS-PEFT variant that uses Adam-style updates for $\tau$. These results provide direct empirical evidence that adaptive momentum estimation is crucial for stabilizing the learning of token-level thresholds in practice.
A.3 Module-wise average sparsity of LLaMA3.1-8B on the commonsense reasoning task
Table 5 reports the detailed module-wise sparsity of TS-LoRA, TS-DoRA, and TS-AdaLoRA on LLaMA3.1-8B for commonsense reasoning. For each transformer layer (indexed from 0 to 31), we measure the token-level sparsity of the PEFT updates in the query, key, and value projections (q_proj, k_proj, and v_proj), and then compute their average (“avg” column). The last row summarizes the mean sparsity over all layers for each method. A dash (“–”) indicates that the corresponding module is not assigned any rank (i.e., effectively pruned by the underlying PEFT scheme).
From the table, we observe that TS-LoRA and TS-AdaLoRA exhibit relatively smooth sparsity profiles across layers: their per-layer averages mostly stay within the 50%–70% range, and no single layer dominates the overall sparsity pattern. This matches the moderate standard deviations reported in the main text (9.8 for TS-LoRA and 11.4 for TS-AdaLoRA), suggesting that these two methods tend to distribute token-level updates more evenly over depth. In contrast, TS-DoRA shows a much more polarized pattern: several layers (e.g., layers 0–2, 5–7, 11, 17, 24, 29) have near-100% sparsity in one or more projections, while a few layers (such as 8, 9, 14, 22) maintain comparatively low sparsity. This leads to a much larger standard deviation (41.3), indicating that TS-DoRA concentrates effective updates into a smaller subset of layers. These fine-grained statistics support our observation in the main paper that the token-level sparsity learned by TS-DoRA can serve as a strong indicator of module importance and motivates the subsequent experiments that select only the low-sparsity (high-importance) modules for fine-tuning.
TS-LoRA TS-DoRA TS-AdaLoRA
Layer q_proj k_proj v_proj avg q_proj k_proj v_proj avg q_proj k_proj v_proj avg
0 65.6 77.6 0.0 47.7 100.0 0.0 0.0 33.3 - - 76.9 76.9
1 59.7 65.5 54.8 60.0 100.0 0.0 65.0 55.0 62.3 84.5 68.4 71.8
2 62.8 64.9 51.5 59.7 100.0 100.0 57.2 85.7 57.3 75.0 81.7 71.3
3 60.1 64.4 56.3 60.2 99.8 100.0 37.1 79.0 67.6 68.7 81.9 72.7
4 63.7 67.6 56.6 62.6 24.6 0.0 83.0 35.9 48.2 70.2 75.8 64.7
5 57.9 62.6 54.8 58.4 0.0 0.0 30.9 10.3 49.6 86.4 80.9 72.3
6 62.4 60.6 48.4 57.1 100.0 0.0 62.9 54.3 70.7 84.8 66.1 73.9
7 61.3 65.9 51.2 59.5 0.0 0.0 82.8 27.6 59.8 86.6 77.2 74.5
8 60.8 65.8 50.4 59.0 70.5 57.1 60.2 62.6 54.0 93.6 60.7 69.4
9 65.9 51.1 53.1 56.7 0.0 11.6 47.4 19.7 52.0 72.4 61.2 61.9
10 57.7 58.7 52.8 56.4 0.0 0.0 41.3 13.8 51.3 82.5 64.2 66.0
11 58.5 57.5 42.3 52.8 91.7 77.2 76.1 81.7 54.2 86.6 65.0 68.6
12 51.9 68.2 49.9 56.7 0.0 0.0 2.1 0.7 50.7 87.9 56.0 64.9
13 58.9 68.3 51.2 59.5 90.2 90.0 35.1 71.7 61.3 73.4 56.1 63.6
14 57.1 60.6 48.9 55.6 0.0 82.6 44.6 42.4 62.0 68.6 44.8 58.5
15 63.2 63.5 51.0 59.3 81.2 0.0 35.2 38.8 66.3 69.0 46.0 60.4
16 58.5 69.9 55.9 61.4 99.8 100.0 3.6 67.8 59.2 86.9 48.8 65.0
17 62.9 70.8 54.7 62.8 99.3 100.0 30.9 76.7 61.6 74.0 53.6 63.1
18 67.3 75.9 61.3 68.2 99.5 70.0 47.0 72.2 62.6 94.7 62.5 73.3
19 65.6 64.4 54.0 61.3 98.0 0.0 78.5 58.8 66.9 80.5 62.7 70.0
20 67.8 66.0 47.5 60.4 97.6 0.0 0.8 32.8 65.5 85.8 54.7 68.6
21 58.2 63.5 44.2 55.3 99.4 100.0 42.4 80.6 61.0 69.1 45.4 58.5
22 62.8 57.6 58.2 59.5 99.8 98.0 79.5 92.4 64.5 70.8 59.7 65.0
23 70.0 71.7 55.6 65.8 99.5 2.5 14.0 38.7 69.6 71.9 57.7 66.4
24 72.3 63.0 53.5 62.9 0.0 0.3 46.0 15.4 68.8 73.7 67.2 69.9
25 69.8 63.2 56.3 63.1 0.2 0.0 4.7 1.6 58.6 65.8 76.9 67.1
26 65.1 70.2 56.4 63.9 9.4 94.2 56.5 53.4 64.4 74.5 67.7 68.8
27 68.9 67.2 68.2 68.1 99.8 95.1 45.1 80.0 62.7 85.8 78.3 75.6
28 57.8 58.9 61.5 59.4 17.9 99.9 94.9 70.9 63.8 76.2 86.1 75.4
29 71.6 78.7 49.3 66.5 0.0 100.0 11.4 37.1 63.9 74.4 79.7 72.7
30 59.5 59.6 57.2 58.8 0.0 100.0 39.2 46.4 67.8 68.6 72.9 69.8
31 64.4 67.3 46.1 59.3 100.0 0.0 1.2 33.7 69.3 71.0 57.7 66.0
avg 64.8 67.4 53.3 60.6 47.7 43.8 61.2 77.9 67.6
Table 5: Module-wise average sparsity (%) of LLaMA3.1-8B on commonsense reasoning. “-” denotes rank 0.
A.4 Hyperparameters
Across all experiments, we tune the sparsity-related hyperparameters s and $\lambda$ separately from the other training configurations. For CSR and VIT, we fix the scaling factor at s = $4\times 10^{-5}$ and select $\lambda$ from the interval $10^{-7},10^{-3}]$ based on development performance. For NLU tasks, we instead set $s=1\times 10^{-4}$ and choose $\lambda$ from $10^{-8},10^{-3}]$. The remaining hyperparameters, including rank, scaling coefficient, dropout rate, optimizer, learning rate, scheduler, batch size, number of epochs, and target modules, are summarized in Table 6.
Task Data Rank Scale Dropout Optimizer LR Scheduler Batch Epochs Target
CSR ALL 32 0.5 0.05 AdamW 1e-4 Linear 16 3 Q,K,V,Up,Down
VIT ALL 128 0.5 0.05 AdamW 2e-4 Cosine 16 1 Q,K,V,O,Up,Down,Gate
NLU CoLA 4 0.5 0.1 AdamW 4e-4 Linear 32 40 Q,K,V,O,I
NLU MNLI 4 0.5 0.15 AdamW 2.5e-4 Linear 32 10 Q,K,V,O,I
NLU QNLI 4 0.5 0.1 AdamW 2e-4 Linear 32 10 Q,K,V,O,I
NLU RTE 4 0.5 0.2 AdamW 1e-4 Linear 32 80 Q,K,V,O,I
NLU MRPC 4 0.5 0 AdamW 1e-4 Linear 32 60 Q,K,V,O,I
NLU QQP 4 0.5 0.15 AdamW 4e-4 Linear 32 10 Q,K,V,O,I
NLU SST-2 4 0.5 0 AdamW 4e-4 Linear 32 48 Q,K,V,O,I
NLU STS-B 4 0.5 0.2 AdamW 5e-4 Linear 32 45 Q,K,V,O,I
Table 6: Hyperparameter configuration for our experiments.

### v2

@@
-TS-PEFT: Token-Selective Parameter-Efficient Fine-Tuning with Learnable Threshold Gating
+TS-PEFT: Unveiling Token-Level Redundancy in Parameter-Efficient Fine-Tuning
 Abstract
-In the field of large models (LMs) for natural language processing (NLP) and computer vision (CV), Parameter-Efficient Fine-Tuning (PEFT) has emerged as a resource-efficient method that modifies a limited number of parameters while keeping the pretrained weights fixed.
-This paper investigates the traditional PEFT approach, which applies modifications to all position indices, and questions its necessity.
-We introduce a new paradigm called Token-Selective PEFT (TS-PEFT), in which a function S selectively applies PEFT modifications to a subset of position indices, potentially enhancing performance on downstream tasks.
-Our experimental results reveal that the indiscriminate application of PEFT to all indices is not only superfluous, but may also be counterproductive.
-This study offers a fresh perspective on PEFT, advocating for a more targeted approach to modifications and providing a framework for future research to optimize the fine-tuning process for large models.
+Current Parameter-Efficient Fine-Tuning (PEFT) methods typically operate under an implicit assumption: once a target module is selected, every token passing through it contributes equally to the downstream task and requires a parameter update.
+In this paper, we challenge this convention and unveil a pervasive token-level redundancy in the fine-tuning of large models.
+We propose TS-PEFT, a theoretically grounded framework utilizing proximal optimization to dynamically identify and skip redundant token updates during training.
+Our extensive experiments across Natural Language Understanding, Commonsense Reasoning, and Visual Instruction Tuning demonstrate that indiscriminately updating all tokens is not only computationally superfluous but often introduces optimization noise.
+Strikingly, by discarding 40%-60% of token updates, TS-PEFT consistently matches or surpasses the performance of dense baselines (e.g., LoRA, DoRA).
+Furthermore, we provide an in-depth analysis revealing that the learned token-level sparsity serves as a superior indicator of module importance compared to traditional weight norms, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
 1 Introduction
-Large models (LMs) have achieved significant progress in natural language processing (NLP) yao2024survey, computer vision (CV) croitoru2023diffusion; liu2024sora, and many other fields.
-However, due to their large parameter scales, directly fine-tuning all parameters is often impractical in terms of computational and storage overhead.
-To address this issue, parameter-efficient fine-tuning (PEFT) has gradually become a mainstream paradigm for adapting pretrained large models: by freezing the base model network and introducing only a small number of trainable parameters, downstream customization can be accomplished efficiently hu2023llm; hanparameter.
-Various PEFT methods have demonstrated strong performance across a wide range of downstream tasks.
+Large models (LMs) have demonstrated remarkable capabilities across Natural Language Processing (NLP) yao2024survey and Computer Vision (CV) croitoru2023diffusion; liu2024sora.
+To adapt these massive models to downstream tasks efficiently, Parameter-Efficient Fine-Tuning (PEFT) has become the de facto paradigmhu2023llm; hanparameter.
 Among them, prompt-based approaches (such as Prefix-Tuning li2021prefix and Prompt-Tuning lester2021power) adjust model behavior by learning additional input tokens, while low-rank adaptation (LoRA) hu2022lora and its variants (e.g., AdaLoRA zhang2023adaptive, DoRA liu2024dora, VeRA kopiczkovera) update internal representations by injecting a small number of trainable parameters or low-rank components into the base model.
 These approaches significantly reduce the number of trainable parameters while typically maintaining performance comparable to, or even approaching, full-parameter fine-tuning.
-Most existing PEFT methods, however, implicitly rely on a key assumption: once a layer is selected as the target module for PEFT, all token positions passing through this module should receive a trainable modification.
+However, despite their distinct architectural designs, these methods share a common, implicit assumption: the ”Dense Update” assumption.
+That is, once a module (e.g., a self-attention layer) is targeted for fine-tuning, the learnable modification is indiscriminately applied to every token position in the input sequence.
 Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is typically written as:
 $h_{i}=W_{0}x_{i}+\operatorname{M}(x_{i}),$ (1)
-where $\operatorname{M}(\cdot)$ is a PEFT module such as LoRA. This “update-all-positions” design ignores the fact that different token positions contribute unequally to the downstream objective.
+where $\operatorname{M}(\cdot)$ is a PEFT module such as LoRA.
 Figure 1: Comparison between standard PEFT and our TS-PEFT framework.
-In practice, we observe that, for already strong pretrained models, the deviation induced by PEFT is extremely small on many token positions, while only a small subset of positions exhibit substantial and task-relevant changes.
-In other words, most PEFT updates are effectively redundant at the token level.
-Uniformly applying modifications to all positions not only injects unnecessary noise and potential interference, but also increases computation, especially when multiple PEFT modules are stacked.
-This motivates a natural question: can PEFT be made selective over token positions instead of blindly updating every one of them?
-Our empirical findings suggest a negative answer.
-We propose a simple yet effective framework, Token-Selective PEFT (TS-PEFT), which enables token-level selective parameter-efficient fine-tuning.
-Instead of always applying $\operatorname{M}(x_{i})$ to every position, TS-PEFT introduces a binary gating function $S$ at each token position that decides whether to add the PEFT update or to keep the base model output unchanged.
-Intuitively, positions whose PEFT updates are small relative to the frozen output remain untouched, while positions with larger, more impactful updates retain their PEFT modifications, yielding a token-level sparse pattern of PEFT updates.
-In the experimental section, we conduct systematic evaluations on representative models including LLaMA-3-8B, LLaVA-1.5-7B, and DeBERTaV3-base hedebertav3, covering commonsense reasoning, visual instruction tuning, and natural language understanding.
-The results show that, on top of strong PEFT baselines such as LoRA, DoRA, AdaLoRA, and Adapter, TS-PEFT consistently matches or surpasses their performance while applying updates to only about 40–60% of token positions.
-Moreover, the learned token-level sparsity patterns can be aggregated into module-level importance indicators, which enable selecting fewer yet more critical tuning modules that even outperform full-module fine-tuning under comparable parameter budgets.
+In this paper, we challenge this convention and investigate a fundamental scientific question: Is it necessary to apply fine-tuning updates to all token positions?
+Given the strong representational power of pre-trained backbones, we hypothesize that for a specific downstream sample, only a subset of tokens requires task-specific adaptation, while others retain sufficient information from the pre-trained weights.
+Under this hypothesis, the standard dense update strategy is not only computationally superfluous but may also introduce optimization noise and interference, potentially hindering model performance.
+Recent works like CFT ruan2025enhancing have demonstrated that masking out non-critical tokens during training improves reasoning.
+However, CFT treats redundancy as a static data filtering problem, requiring costly offline counterfactual perturbations to generate fixed masks.
+In contrast, to systematically verify this hypothesis and quantify the potential redundancy, we propose Token-Selective PEFT (TS-PEFT).
+Unlike prior approaches, TS-PEFT serves as a learnable probe that dynamically decides whether to update a token or keep its representation frozen.
+We formulate this selection process as a sparsity-regularized optimization problem, employing proximal optimization and an Adam-style adaptive thresholding mechanism to ensure stable, end-to-end training of the gating decisions.
+Our empirical findings unveil a pervasive token-level redundancy in standard PEFT.
+Across diverse benchmarks including Commonsense Reasoning, Visual Instruction Tuning, and Natural Language Understanding, we observe that discarding 40%-60% of token updates does not degrade performance.
+Strikingly, TS-PEFT often outperforms its dense counterparts (e.g., LoRA, DoRA) by explicitly filtering out redundant updates.
+While current hardware constraints on sparse operations limit immediate wall-clock speedups, our work establishes a theoretical computational bound, demonstrating that nearly half of the fine-tuning computation in current paradigms is effectively unnecessary.
+Furthermore, our analysis exposes a profound connection between token sparsity and module importance.
+We discover that the learned sparsity pattern is not random; modules that exhibit low sparsity (i.e., require frequent updates) are significantly more critical to the downstream task than those with high sparsity.
+This insight allows token-level sparsity to serve as a superior indicator for module selection compared to traditional weight-norm-based metrics.
 In summary, our contributions are as follows:
-  • We point out and examine a neglected aspect of standard PEFT: once a layer is chosen as a target module, all token positions are updated uniformly, even though only a small fraction of positions contribute meaningfully to the downstream objective.
-Our empirical analysis of token-level sparsity shows that many PEFT updates induce negligible deviation from the frozen backbone, revealing substantial redundancy along the sequence dimension and motivating selective updates.
-  • We propose TS-PEFT, which introduces a token-level binary gating function on top of existing PEFT modules.
-By formulating the learning of layer-wise thresholds as a sparsity-regularized proximal optimization problem, deriving a practical approximate gradient, and applying Adam-style moment estimation, TS-PEFT enables stable end-to-end training of token-level gates jointly with PEFT parameters.
-The framework is agnostic to the specific low-rank parameterization and can be used as a plug-in module for a variety of PEFT methods.
-  • On three representative benchmarks, TS-PEFT consistently matches or surpasses strong PEFT baselines, while updating only 40%-60% token positions.
-Moreover, the learned token-level sparsity patterns can be aggregated into module-wise importance scores, enabling selective tuning of fewer but more critical modules that can outperform full-module fine-tuning under comparable parameter budgets.
+  • Unveiling Redundancy: We identify and quantify that nearly 40%-60% of token updates in standard PEFT are redundant.
+We empirically demonstrate that indiscriminately updating these tokens is suboptimal and introduces optimization noise.
+  • Methodological Framework: We propose TS-PEFT, a theoretically grounded framework that employs proximal optimization to dynamically learn token-selective policies in an end-to-end manner, bypassing the need for costly offline heuristics.
+  • Mechanism Insight: We reveal that token-level sparsity provides a strong signal for module importance, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
 2 Related Work
@@
 Building on this idea, AdaLoRA zhang2023adaptive adaptively redistributes the effective rank across different layers to better meet task-specific requirements, DoRA liu2024dora refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA kopiczkovera leverages shared low-rank bases to further reduce per-task parameter overhead.
-However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update—that is, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension.
+However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update.
+That is to say, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension.
 This may introduce unnecessary redundant modifications and fails to explicitly characterize or exploit potential token-level redundancy.
@@
 tan2024sparsity further introduce sparsity-guided techniques as a tool for interpreting the behavior of LLMs.
-However, most existing studies focus on parameter-wise sparsity, i.e., deciding which weights or modules to activate or prune.
-In contrast, our work explores token-level sparsity in the context of PEFT: instead of sparsifying parameters, we selectively decide which token positions should receive PEFT updates, revealing and leveraging redundancy along the sequence dimension.
+We observe that these methods exploit redundancy at different levels.
+MoE-based approaches typically target modular-level sparsity, whereas pruning focuses on parameter-level sparsity.
+In contrast, our work explores token-level sparsity within the context of PEFT, introducing a gating mechanism to selectively skip updates for tokens requiring no task-specific adaptation.
+Therefore, our method is independent of these architectures and can be theoretically plugged in as a detachable module on top of LoRA-MoE, further eliminating redundant computations.
 3 Method
@@
 In practice, this Adam-style update yields stable convergence of the thresholds under noisy gradients while preserving the desired sparsity level.
+Optimization Stability.
+We empirically observe that the adaptive moment estimation (Eq.
+(16)-Eq.
+(20)) is critical for convergence.
+Ablation studies in Appendix A.2 demonstrate that removing this component leads to severe threshold oscillation and training divergence, resulting in significant performance degradation on sensitive tasks like BoolQ.
 Hyperparams PIQA BoolQ HellaSwag WinoGrande SIQA OBQA ARC-e ARC-c Avg
@@
 The results show that TS-Adapter performs no worse than the original Adapter, suggesting that TS-PEFT is not bound to LoRA-style weight structures, but can function as a general token-level gating layer that can be stacked on diverse PEFT modules.
-TS-PEFT introduces token-level sparsity as a logical mask.
-Therefore, the current implementation does not provide wall-clock speedups.
-However, the theoretical computation reduction is significant.
-With specialized kernels support for sparse matrix additions in the future, TS-PEFT has the potential to translate this logical sparsity into physical latency reduction.
-4.3 In-depth Analysis
+Although TS-PEFT achieves a theoretical computation reduction of approximately 40%-60% in FLOPs within the adapter modules, current deep learning accelerators and libraries remain heavily optimized for dense matrix operations.
+Consequently, the imposed token-level sparsity currently serves as a theoretical computational bound and a probe for quantifying redundancy, rather than an immediate mechanism for wall-clock speedup.
+Nevertheless, with the emerging support for sparse tensor cores in next-generation hardware, TS-PEFT establishes a blueprint for translating this identified redundancy into tangible physical efficiency.
+4.3 Analysis: Token Sparsity Reveals Model Adaptation Mechanism
 Figure 2: Performance of LLaMA3.1-8B on CSR benchmarks as the percentage of selected modules varies for each PEFT method.
+Note the performance drops (e.g., DoRA at 60%) when redundant, high-sparsity modules are forced to update, indicating noise injection.
 PIQA BoolQ HellaSwag WinoGrande SIQA OBQA ARC-e ARC-c Avg
@@
 half_rank_50% 88.1 75.4 95.2 86.3 81.2 85.2 89.6 78.7 84.9
-Table 4: The performance of LLaMA3.1-8B on the commonsense reasoning task using various selection methods.
-Necessity of Adaptive Moment Estimation.
-To verify the effect of the adaptive momentum estimation module, we conduct ablation experiments by removing this component.
-In this case, the threshold $\tau$ exhibits noticeable oscillation and gradually diverges, and the training loss also fails to converge.
-On representative CSR tasks such as BoolQ, this instability leads to a clear degradation in dev performance compared with the full TS-PEFT variant (see Appendix A.2 for detailed curves and numerical results).
-These observations indicate that adaptive momentum estimation is essential for stabilizing the threshold update process in practice.
-Module-wise Sparsity Distribution.
+Table 4: Performance of LLaMA3.1-8B on the commonsense reasoning task using various selection methods.
+Sparsity Patterns Reflect Architecture Bias.
 In LLaMA3.1-8B, there are 32 layers in total.
@@
 This indicates that, in our setting, the token-level sparsity patterns learned by TS-DoRA provide a more reliable signal of module importance than simple norm statistics or random selection.
-Effect of the Number of Target Modules.
-After using sparsity as the criterion for module importance ranking, we further examine the effect of varying the proportion of target modules while keeping the ranking fixed: we select different percentages of modules in ascending order of sparsity and fine-tune them.
-The results are shown in Figure 2.
-For AdaLoRA, performance remains largely unaffected as the proportion of target modules increases.
-However, for the others, “more is better” does not hold: DoRA achieves its best performance when using only 50% of the modules, LoRA reaches its optimal performance at 80%, and even using only 20% of the modules yields results close to the optimum.
-In addition, LoRA and DoRA exhibit clear performance cliffs at certain points: for DoRA, performance drops by about 6.6% when increasing the target module proportion from 50% to 60%; for LoRA, increasing the proportion from 90% to 100% results in a drop of about 1.3%.
-This phenomenon indicates that jointly training modules with higher sparsity may negatively interfere with overall performance.
-In contrast, AdaLoRA is less sensitive to such “low-importance modules,” as its dynamic rank allocation redistributes capacity across modules.
+Redundant Updates Introduce Optimization Noise.
+To further validate that the identified high-sparsity modules are indeed redundant and potentially harmful, we conducted a cumulative selection analysis.
+We expanded the set of target modules in ascending order of sparsity (i.e., adding modules from “critical” to “redundant”).
+As illustrated in Figure 2, the conventional wisdom that ”more trainable parameters imply better performance” does not hold.
+For DoRA, performance peaks when only 50% of the modules are fine-tuned, while LoRA reaches its optimum at 80%.
+Remarkably, even utilizing just 20% of the modules yields results comparable to the optimal settings
+Crucially, we observe distinct “performance cliffs” when redundant modules are forced to update.
+For instance, DoRA suffers a sharp performance drop of 6.6% when the target proportion increases from 50% to 60%.
+Similarly, LoRA experiences a 1.3% degradation when expanding from 90% to 100%.
+This phenomenon provides compelling evidence that updating modules characterized by high sparsity injects negative interference rather than meaningful adaptation.
+In contrast, AdaLoRA exhibits greater stability against these low-importance modules, likely due to its dynamic rank allocation mechanism which automatically limits the capacity assigned to redundant positions.
 5 Conclusion
@@
 We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 40–60% of token positions, yet consistently matches or surpasses strong baselines.
-Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a basis for future work on more efficient sparse fine-tuning and inference.
+Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a theoretical basis for future hardware-software co-design targeting efficient sparse fine-tuning and inference.
 Appendix A Appendix

### v3

@@
 Current Parameter-Efficient Fine-Tuning (PEFT) methods typically operate under an implicit assumption: Once a target module is selected, every token passing through it contributes equally to the downstream task and requires a parameter update.
-In this paper, we challenge this convention and unveil a pervasive token-level redundancy in the fine-tuning of large models.
-We propose TS-PEFT, a theoretically grounded framework utilizing proximal optimization to dynamically identify and skip redundant token updates during training.
-Our extensive experiments across Natural Language Understanding, Commonsense Reasoning, and Visual Instruction Tuning demonstrate that indiscriminately updating all tokens is not only computationally superfluous but often introduces optimization noise.
-Strikingly, by discarding 40%-60% of token updates, TS-PEFT consistently matches or surpasses the performance of dense baselines (e.g., LoRA, DoRA).
-Furthermore, we provide an in-depth analysis revealing that the learned token-level sparsity serves as a superior indicator of module importance compared to traditional weight norms, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
+In this paper, we challenge this convention by revealing a pervasive token-level redundancy in the fine-tuning of large models (LMs).
+We propose TS-PEFT, a theoretical framework utilizing proximal optimization that acts as a dynamic probe to identify token-level redundancy during the fine-tuning process.
+Extensive experiments demonstrate that indiscriminately updating all tokens is not only computationally superfluous but often introduces optimization noise.
+Surprisingly, by discarding 30%-70% of token updates, TS-PEFT consistently matches or exceeds the performance of dense baselines such as LoRA, DoRA.
+Our in-depth analysis shows that the learned token-level sparsity is a superior indicator of module importance compared to traditional weight criteria, providing a novel data-driven perspective on the intrinsic adaptation mechanism of LMs.
 1 Introduction
-Large models (LMs) have demonstrated remarkable capabilities across Natural Language Processing (NLP) yao2024survey and Computer Vision (CV) croitoru2023diffusion; liu2024sora.
-To adapt these massive models to downstream tasks efficiently, Parameter-Efficient Fine-Tuning (PEFT) has become the de facto paradigmhu2023llm; hanparameter.
-Among them, prompt-based approaches (such as Prefix-Tuning li2021prefix and Prompt-Tuning lester2021power) adjust model behavior by learning additional input tokens, while low-rank adaptation (LoRA) hu2022lora and its variants (e.g., AdaLoRA zhang2023adaptive, DoRA liu2024dora, VeRA kopiczkovera) update internal representations by injecting a small number of trainable parameters or low-rank components into the base model.
-These approaches significantly reduce the number of trainable parameters while typically maintaining performance comparable to, or even approaching, full-parameter fine-tuning.
-However, despite their distinct architectural designs, these methods share a common, implicit assumption: the ”Dense Update” assumption.
-That is, once a module (e.g., a self-attention layer) is targeted for fine-tuning, the learnable modification is indiscriminately applied to every token position in the input sequence.
-Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is typically written as:
+Large models (LMs) are demonstrating astonishing capabilities in Natural Language Processing (NLP) Yao et al.
+(2024); Guo et al.
+(2025) and Computer Vision (CV) Croitoru et al.
+(2023); Liu et al.
+(2024d).
+To effectively use these large-scale models in downstream tasks, the dominant paradigm is Parameter-Efficient Fine-Tuning (PEFT)Hu et al.
+(2023); Han et al.
+(2024).
+In these approaches, prompt-based methods such as Prefix-Tuning Li and Liang (2021) and Prompt-Tuning Lester et al.
+(2021) train additional input tokens to change model behavior and low-rank adaptation (LoRA) Hu et al.
+(2022) and its variants Zhang et al.
+(2023); Liu et al.
+(2024c, c) change internal representations by injecting a small number of trainable parameters or corridors, i.e.
+low-rank components, into the base model.
+These methods effectively decrease trainable parameters while maintaining similar, or comparable, performance to full parameter fine-tuning.
+However, despite their distinct architectural designs, these methods share a common and implicit assumption that they all employ dense updates.
+That is, once a module like a self-attention layer is used for fine-tuning, the learnable modification is applied indiscriminately to every token position in the input sequence.
+Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is usually written as:
 $h_{i}=W_{0}x_{i}+\operatorname{M}(x_{i}),$ (1)
@@
 In this paper, we challenge this convention and investigate a fundamental scientific question: Is it necessary to apply fine-tuning updates to all token positions?
-Given the strong representational power of pre-trained backbones, we hypothesize that for a specific downstream sample, only a subset of tokens requires task-specific adaptation, while others retain sufficient information from the pre-trained weights.
-Under this hypothesis, the standard dense update strategy is not only computationally superfluous but may also introduce optimization noise and interference, potentially hindering model performance.
-Recent works like CFT ruan2025enhancing have demonstrated that masking out non-critical tokens during training improves reasoning.
-However, CFT treats redundancy as a static data filtering problem, requiring costly offline counterfactual perturbations to generate fixed masks.
-In contrast, to systematically verify this hypothesis and quantify the potential redundancy, we propose Token-Selective PEFT (TS-PEFT).
+Given the strong representational power of pretrained base models, we assume that for a specific downstream sample, only a subset of tokens requires task-specific adaptation, while others retain sufficient information from the pretrained weights.
+Under this assumption, the standard dense update strategy is suboptimal, primarily because it introduces optimization noise that forces the model to fit gradients on irrelevant tokens, distorting the well-learned features.
+Recent works like CFT Ruan et al.
+(2025) have demonstrated that masking non-critical tokens during training can improve reasoning.
+This finding corroborates our intuition that, just as imposing a uniform penalty on all tokens in pretraining may hurt generalization, updating all tokens indiscriminately in PEFT similarly introduces redundancy and noise.
+To systematically verify this assumption and quantify the potential redundancy, we propose Token-Selective PEFT (TS-PEFT).
 Unlike prior approaches, TS-PEFT serves as a learnable probe that dynamically decides whether to update a token or keep its representation frozen.
 We formulate this selection process as a sparsity-regularized optimization problem, employing proximal optimization and an Adam-style adaptive thresholding mechanism to ensure stable, end-to-end training of the gating decisions.
-Our empirical findings unveil a pervasive token-level redundancy in standard PEFT.
-Across diverse benchmarks including Commonsense Reasoning, Visual Instruction Tuning, and Natural Language Understanding, we observe that discarding 40%-60% of token updates does not degrade performance.
-Strikingly, TS-PEFT often outperforms its dense counterparts (e.g., LoRA, DoRA) by explicitly filtering out redundant updates.
-While current hardware constraints on sparse operations limit immediate wall-clock speedups, our work establishes a theoretical computational bound, demonstrating that nearly half of the fine-tuning computation in current paradigms is effectively unnecessary.
+Our empirical findings unveil a pervasive token-level redundancy.
+Across many benchmarks, we observe that nearly 30%-70% of token updates are actually redundant.
+Crucially, we prove that indiscriminately updating these tokens is detrimental.
+As evidenced by our analysis, forcing updates on these redundant positions leads to a “performance cliff”, confirming that redundancy in PEFT acts as harmful noise.
+By filtering out this noise, TS-PEFT consistently matches or exceeds the performance of dense baselines (e.g., LoRA, DoRA) while utilizing significantly fewer active parameters.
 Furthermore, our analysis exposes a profound connection between token sparsity and module importance.
-We discover that the learned sparsity pattern is not random; modules that exhibit low sparsity (i.e., require frequent updates) are significantly more critical to the downstream task than those with high sparsity.
+We discover that the learned sparsity pattern is not random.
+The modules that exhibit low sparsity, i.e., requiring frequent updates, are significantly more important to the downstream task than those with high sparsity.
 This insight allows token-level sparsity to serve as a superior indicator for module selection compared to traditional weight-norm-based metrics.
 In summary, our contributions are as follows:
-  • Unveiling Redundancy: We identify and quantify that nearly 40%-60% of token updates in standard PEFT are redundant.
+  • Unveiling Redundancy: We identify and quantify that in standard PEFT, nearly 30%-70% of token updates are redundant.
 We empirically demonstrate that indiscriminately updating these tokens is suboptimal and introduces optimization noise.
-  • Methodological Framework: We propose TS-PEFT, a theoretically grounded framework that employs proximal optimization to dynamically learn token-selective policies in an end-to-end manner, bypassing the need for costly offline heuristics.
-  • Mechanism Insight: We reveal that token-level sparsity provides a strong signal for module importance, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
+  • Methodological Framework: We propose TS-PEFT, a theoretical framework that employs proximal optimization to dynamically learn token-selective strategies in an end-to-end manner.
+  • Mechanism Insight: We reveal that token-level sparsity provides a strong signal for module importance, offering a novel data-driven perspective on the intrinsic adaptation mechanism of LMs.
 2 Related Work
 Low-Rank Adaptation.
-Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while preserving downstream performance.
-LoRA hu2022lora introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models; after training, these matrices can be merged back into the base model without incurring additional inference cost.
-Building on this idea, AdaLoRA zhang2023adaptive adaptively redistributes the effective rank across different layers to better meet task-specific requirements, DoRA liu2024dora refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA kopiczkovera leverages shared low-rank bases to further reduce per-task parameter overhead.
+Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while maintaining downstream performance.
+LoRA Hu et al.
+(2022) introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models.
+After training, these matrices can be merged back into the base model without incurring additional inference cost.
+Based on this concept, AdaLoRA Zhang et al.
+(2023) adaptively redistributes the effective rank across different layers to better meet task-specific requirements.
+DoRA Liu et al.
+(2024c) refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA Kopiczko et al.
+(2024) utilizes shared low-rank bases to further reduce the parameter overhead for each task.
 However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update.
@@
 Sparsity.
-Sparsity-guided methods have been widely used to improve the efficiency and interpretability of large language models.
+The methods of sparsity guidance have been widely used to improve the efficiency and interpretability of large language models (LLMs).
 By reducing the number of active parameters and concentrating computation on core components, these approaches not only make models more efficient but also offer clearer insights into their decision-making processes.
-Mixture-of-experts (MoE) models, exemplified by Switch Transformers fedus2022switch and Mixtral 8x7B jiang2024mixtral, employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters, and several recent works have combined MoE architectures with LoRA-style adapters to further enhance performance chen2024llava; liu2024moe; zadouripushing .
-Another typical use of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets ma2023llm; xubesa; xiasheared.
-tan2024sparsity further introduce sparsity-guided techniques as a tool for interpreting the behavior of LLMs.
-We observe that these methods exploit redundancy at different levels.
+Mixture-of-experts (MoE) models, such as Switch Transformers Fedus et al.
+(2022) and Mixtral 8x7B Jiang et al.
+(2024), employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters.
+Recently, several works have combined MoE architectures with LoRA-style adapters to further enhance performance Chen et al.
+(2024); Liu et al.
+(2024b); Zadouri et al.
+(2024).
+Another typical application of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets Ma et al.
+(2023); Xu et al.
+(2024); Xia et al.
+(2024).
+Tan et al.
+(2024) further introduces sparsity-guided techniques as a tool for interpreting the behavior of LLMs.
+We observe that these methods all utilize redundancy at different levels.
 MoE-based approaches typically target modular-level sparsity, whereas pruning focuses on parameter-level sparsity.
 In contrast, our work explores token-level sparsity within the context of PEFT, introducing a gating mechanism to selectively skip updates for tokens requiring no task-specific adaptation.
-Therefore, our method is independent of these architectures and can be theoretically plugged in as a detachable module on top of LoRA-MoE, further eliminating redundant computations.
+Therefore, our approach is not constrained by these architectures, and theoretically it can be directly embedded as a detachable module into all LoRA variants, including LoRA-MoE, thereby further eliminating redundant computations.
 3 Method
-This section presents the overall theoretical framework and optimization process of TS-PEFT.
+This section introduces the overall theoretical framework and optimization process of TS-PEFT.
 We first formalize the LoRA-style PEFT procedure and introduce a token-level control function $\operatorname{S}$ to determine whether a PEFT update should be applied at each position.
@@
 Next, to address the non-differentiability of $\operatorname{S}$ with respect to $\tau$ and the issue that gradients are nearly zero at most positions, we derive a computable approximate gradient.
-Finally, we combine this approximate gradient with momentum and adaptive learning-rate mechanisms so that it can be seamlessly integrated into the Adam optimizer, and we provide the overall algorithmic description of the TS-PEFT training process.
+Finally, we combine this approximate gradient with momentum and adaptive learning rate mechanisms so that it can be seamlessly integrated into the Adam optimizer, and we provide the overall algorithmic description of the TS-PEFT training process.
 3.1 Token-level Selective PEFT Formulation
@@
 $h_{i}=\begin{cases}W_{0}x_{i}&\text{if }\operatorname{S}(W_{0}x_{i},\operatorname{M}(x_{i}))=0,\\ W_{0}x_{i}+\operatorname{M}(x_{i})&\text{if }\operatorname{S}(W_{0}x_{i},\operatorname{M}(x_{i}))=1.\end{cases}$ (2)
-Here, $\operatorname{S}(\cdot)\in\{0,1\}$ serves as a gating variable: when $\operatorname{S}=1$, the update is enabled at that position, and when $\operatorname{S}=0$, the position retains the base model output.
+Here, $\operatorname{S}(\cdot)\in\{0,1\}$ serves as a gating variable.
+When $\operatorname{S}=1$, the update is enabled at that position, while when $\operatorname{S}=0$, the position retains the base model output.
 To instantiate Eq.(2), we focus on the relative magnitude $r_{i}$ of the PEFT update with respect to the base model output at each position $i$: $r_{i}=\|\operatorname{M}(x_{i})\|_{2}/\|W_{0}x_{i}\|_{2}$.
@@
 The relative values enable each layer to adaptively adjust according to its own weight change ratio.
-Intuitively, a larger $r_{i}$ means that the PEFT update induces a stronger perturbation relative to the base model output at that position and is more likely to be task-relevant, whereas a small $r_{i}$ indicates that the modification is minor and can be treated as negligible.
+Intuitively, a larger $r_{i}$ means that the PEFT update induces a stronger perturbation relative to the base model output at that position and is more likely to be relevant to the task, whereas a small $r_{i}$ indicates that the modification is minor and can be treated as negligible.
 Therefore, we introduce a learnable scalar threshold $\tau$ for each target layer and define the control function accordingly:
 $\operatorname{S}(W_{0},\operatorname{M}(x_{i}))=\begin{cases}0&\text{if }r_{i}<\tau,\\ 1&\text{if }r_{i}\geq\tau.\end{cases}$ (3)
-In other words, when the relative magnitude at a given position is not lower than $\tau$, we consider the update at that position worth retaining; otherwise, the update for that position is skipped.
-This results in a token-level sparse update pattern along the sequence dimension within each selected PEFT module, with the threshold $\tau$ directly controlling the overall sparsity level.
-However, this threshold-based step gating introduces two training challenges: first, $\operatorname{S}$ is a non-smooth step function with respect to $\tau$, so standard backpropagation provides almost no useful gradients; second, we want $\operatorname{S}$ to activate PEFT updates only on a subset of token positions while keeping the remaining ones identical to the backbone, which requires explicit control of the overall sparsity level during threshold learning.
+In other words, if the relative magnitude at a certain position is not lower than $\tau$, we consider the update at that position worth retaining.
+Otherwise, the update for that position is skipped.
+This will form a token-level sparse update pattern along the sequence dimension within each selected PEFT module, and the threshold $\tau$ controls the overall sparsity level.
+However, this threshold-based step gating introduces two training challenges: First, $\operatorname{S}$ is a non-smooth step function with respect to $\tau$, so standard backpropagation provides almost no useful gradients.
+Second, we want $\operatorname{S}$ to activate PEFT updates only on a subset of token positions while keeping the remaining ones identical to the base model.
+This requires explicit control of the overall sparsity level during threshold learning.
 The next subsections tackle these two issues via a proximal objective and a computable gradient approximation.
 3.2 Proximal Optimization for the Threshold
-Due to the training difficulties mentioned above, directly using first-order optimization methods cannot effectively update $\tau$.
-Therefore, to ensure that the threshold can be stably updated during training, we formulate its learning process as a proximal optimization problem, allowing it to follow the changes in the original task loss while explicitly favoring a sparser update pattern.
+Due to the aforementioned training difficulties, directly using first-order optimization methods cannot effectively update $\tau$.
+Thus, in order to ensure that the threshold can be updated stably during the training process, we formulate its learning process as a proximal optimization problem, enabling it to follow the changes in the original task loss and explicitly favor a sparser update pattern.
 Specifically, our objective function is defined as follows:
 $\min\mathcal{L}(\tau)=\ell(\tau)+\lambda\sum_{i}\mathbbm{1}(r_{i}\geq\tau),$ (4)
-where $\ell(\tau)$ denotes the loss of the target task, and the second term penalizes the number of excessively activated tokens to encourage a sparser update, $\mathbbm{1}(\cdot)$ is the indicator function, and $\lambda$ serves as a hyperparameter that balances the sparsity level.
-However, since the indicator function is non-differentiable, we adopt proximal optimization techniques and introduce a quadratic regularization term in the threshold update to limit its fluctuation parikh2014proximal.
+where $\ell(\tau)$ denotes the loss of the target task, and the second term penalizes the number of excessively activated tokens to encourage a sparser update.
+$\mathbbm{1}(\cdot)$ is an indicator function, and $\lambda$, as a hyperparameter, is used to balance the sparsity level.
+We employ proximal optimization to derive the update rule for parameter $\tau$ within a stochastic gradient descent (SGD) framework Parikh et al.
+(2014).
 Given the threshold $\tau^{k}$ at the $k$-th iteration, the update $\tau^{k+1}$ can be written as:
-$\underset{\tau}{\mathrm{argmin}}\left[\ell(\tau)+\frac{\partial\ell}{\partial\tau}(\tau-\tau^{k})+\frac{1}{2\alpha_{k}}\left|\tau-\tau^{k}\right|^{2}+\lambda P(\tau)\right],$ (5)
+$\underset{\tau}{\mathrm{argmin}}\left[\ell(\tau^{k})+\frac{\partial\ell}{\partial\tau}(\tau-\tau^{k})+\frac{1}{2\alpha_{k}}\left|\tau-\tau^{k}\right|^{2}+\lambda P(\tau)\right],$ (5)
 where $\alpha_{k}$ is the step size.
@@
 To solve Eq.(5), we need to compute $\partial\ell/\partial\tau$ and $\partial P/\partial\tau$.
-Thus, we first expand the output in Eq.(2) along the hidden dimension.
+Thus, we first expand the output of Eq.(2) along the hidden dimension.
 Let the hidden representation of the $i$-th token at dimension $j\in\{0,\dots,H-1\}$ be denoted as:
@@
 3.3 Gradient Approximation for the Step Function
-In the previous subsection, we obtained the derivative expressions of the loss function and the sparsity penalty term with respect to the threshold $\tau$.
-However, all these derivatives depend on the partial derivative of the step gating $\mathbbm{1}(r_{i}\geq\tau)$ with respect to $\tau$, whose numerical value is almost zero everywhere, making it impossible to directly perform gradient updates.
-Therefore, to enable stable learning of the threshold through backpropagation, this section constructs a computable gradient approximation (based on the perspective of proximal optimization) to replace the non-differentiable derivative of the indicator function.
+In the previous subsection, we obtain the derivative expressions of the loss function and the sparsity penalty term with respect to the threshold $\tau$.
+However, all these derivatives depend on the partial derivative of the step gating $\mathbbm{1}(r_{i}\geq\tau)$ with respect to $\tau$, and its numerical value is almost zero everywhere, making it impossible to directly perform gradient updates.
+Therefore, to enable stable learning of the threshold through backpropagation, this section constructs a computable gradient approximation based on the perspective of proximal optimization to replace the non-differentiable derivative of the indicator function.
 Algorithm 1 Forward and Backward
@@
 10:   Update $\tau^{k+1}$ as in Eq.(20).
-A straightforward approximation is to treat the gradient of the step function as a constant s,s>0-s,s>0, that is:
+A straightforward approximation, inspired by Bengio et al. (2013), is to treat the gradient of the step function as a constant s,s>0-s,s>0, that is:
 $\frac{\partial}{\partial\tau}\mathbbm{1}\left(r_{i}\geq\tau\right)\approx-s.$ (10)
-Substituting this approximation into Eq.(7)-(9) yields the “rough” gradients of the loss term and the sparsity term with respect to the threshold:
+Substituting this approximation value into Eq.(7)-(9) yields the rough gradients of the loss term and the sparsity term with respect to the threshold:
 $\frac{\partial\ell}{\partial\tau}$ $\approx-s\sum_{i}\mu_{i},$ (11)
 $\frac{\partial P}{\partial\tau}$ $\approx-s\sum_{i}\lambda.$ (12)
-However, we observe that Eq.(12) will always push $\tau$ upward, while the direction of Eq.(11) depends on the sign of $\mu_{i}$, making Eq.(10) a rather coarse approximation.
-To avoid unstable oscillation, we further introduce a consistency mask so that only tokens whose PEFT updates “should be activated” contribute to the threshold update.
-Concretely, we only use positions where the current gating decision agrees with their importance estimate ($\mathbbm{1}(\mu_{i}\geq 0)=\mathbbm{1}(r_{i}\geq\tau)$).
-Therefore, we revise Eq.(11) and Eq.(12) as follows:
+Eq.(10) is a rather rough approximation, which causes highly oscillatory updates of $\tau$ and results in poor model performance.
+We believe this is because Eq.(10) should be 0 for certain value domains, but for the update, we set it to s-s.
+Therefore, to limit this effect, we add indicator conditions.
+We modify Eq.(11) and Eq.(12) as follows:
 $\frac{\partial\ell}{\partial\tau}=-s$ $\sum_{i}\mathbbm{1}\left[\mathbbm{1}\left(\mu_{i}\geq 0\right)=\mathbbm{1}\left(r_{i}\geq\tau\right)\right]\mu_{i},$ (13)
 $\frac{\partial P}{\partial\tau}=-s\sum_{i}\mathbbm{1}\left(r_{i}\geq\tau\right)\lambda.$ (14)
-The constraints can be interpreted as follows: whenever $\tau$ is either large enough ($\mu_{i}{\geq}0$, $r_{i}{<}\tau$) or small enough ($\mu_{i}{\leq}0$, $r_{i}{\geq}\tau$) to render the update of $\tau$ unnecessary, Eq.(13) zeros $\partial\ell/\ \partial\tau$; similarly, when $\tau$ is large enough ($r_{i}{<}\tau$, $\lambda$ is a positive constant), Eq.(14) nullifies $\partial P/\ \partial\tau$.
+The constraints can be understood as follows: Whenever $\tau$ is either large enough ($\mu_{i}{\geq}0$, $r_{i}{<}\tau$) or small enough ($\mu_{i}{\leq}0$, $r_{i}{\geq}\tau$) to render the update of $\tau$ unnecessary, Eq.(13) zeros $\partial\ell/\ \partial\tau$.
+Similarly, when $\tau$ is large enough ($r_{i}{<}\tau$, $\lambda$ is a positive constant), Eq.(14) nullifies $\partial P/\ \partial\tau$.
 Combining both parts, the approximate gradient for updating the threshold can be unified as:
@@
 3.4 Adam-style Threshold Update and Overall Training Procedure
-In practice, the approximate gradient $g_{k}$ for the threshold $\tau$ is highly noisy and varies in scale across training steps: the contributions of tokens change significantly between batches, and the magnitude of $g_{k}$ can fluctuate during optimization.
-Directly applying plain gradient descent to $\tau$ therefore tends to cause oscillation and slow or unstable convergence.
-To mitigate this, we update each threshold with an Adam-style rule that combines momentum and adaptive learning rateskinga2015method.
+In fact, for the threshold $\tau$, its approximate gradient $g_{k}$ is highly noisy and varies in scale across training steps.
+The contributions of tokens change significantly between batches, and the magnitude of $g_{k}$ can fluctuate during optimization.
+Directly applying plain gradient descent to $\tau$ often leads to oscillations and can slow down or make the convergence unstable.
+To mitigate this, we adopt a rule similar to Adam to update, which combines momentum and adaptive learning rate Kingma et al.
+(2015).
 At the $k$-th iteration, we use the approximate gradient $g_{k}$ obtained in the previous subsection to update the first and second moments:
@@
 $\hat{v}_{k}$ $\leftarrow v_{k}/(1-\beta_{2}^{k}),$ (19)
-where $\beta_{1}$ controls the contribution of past gradients and $\beta_{2}$ governs the decay rate of squared gradients, while $\hat{m}_{k}$ and $\hat{v}_{k}$ denote the bias-corrected first- and second-moment estimates used in the Adam update.
+where $\beta_{1}$ controls the contribution of past gradients and $\beta_{2}$ governs the decay rate of squared gradients.
+$\hat{m}_{k}$ and $\hat{v}_{k}$ denote the bias-corrected first- and second-moment estimates used in the Adam update.
 Finally, the threshold is updated as:
 […]
-where $s$ is the gradient scaling factor, and $\epsilon$ prevents division by zero.
-In practice, this Adam-style update yields stable convergence of the thresholds under noisy gradients while preserving the desired sparsity level.
-Optimization Stability.
-We empirically observe that the adaptive moment estimation (Eq.
-(16)-Eq.
-(20)) is critical for convergence.
-Ablation studies in Appendix A.2 demonstrate that removing this component leads to severe threshold oscillation and training divergence, resulting in significant performance degradation on sensitive tasks like BoolQ.
+where $s$ is the gradient scaling factor, and $\epsilon$ is a minuscule constant to prevent division by zero.
+We empirically observe that the adaptive moment estimation (Eq.(16)-Eq.(20)) is critical for convergence.
+Ablation studies in Appendix A.4 demonstrate that removing this component leads to severe threshold oscillation and training divergence, resulting in significant performance degradation on sensitive tasks.
 Hyperparams PIQA BoolQ HellaSwag WinoGrande SIQA OBQA ARC-e ARC-c Avg
@@
 Table 1: Evaluation scores of LLaMA3.1-8B on CSR benchmarks; the last row in each block shows the sparsity of TS-PEFT.
-Combining the above components, the training process of TS-PEFT can be summarized as follows: in each training step, during the forward pass, we perform gating based on each token’s relative update magnitude $r_{i}$ and the threshold $\tau$ of the corresponding layer, applying updates only at positions that satisfy the condition while keeping the backbone output unchanged at all other positions.
-Then, standard backpropagation is used to update all PEFT parameters, while the backbone parameters remain frozen throughout.
+Combining the above components, the training process of TS-PEFT can be summarized as follows: In each training step, during the forward propagation phase, we perform gating based on each token’s relative update magnitude $r_{i}$ and the threshold $\tau$ of the corresponding layer, applying updates only at positions that satisfy the condition while keeping the base model output unchanged at all other positions.
+Then, standard backpropagation is used to update all PEFT parameters, while the base model parameters remain frozen throughout.
 On this basis, we compute $\mu_{i}$ to construct the approximate gradient $g_{k}$ for the threshold, and independently update the threshold $\tau$ of each layer using Adam-style first- and second-moment estimates.
@@
 4.1 Experimental Setup
-We evaluate TS-PEFT across three representative scenarios: natural language understanding (NLU), commonsense reasoning (CSR), and visual instruction tuning (VIT).
+We evaluate TS-PEFT in multiple representative scenarios: Natural Language Understanding (NLU), Commonsense Reasoning (CSR), Visual Instruction Tuning (VIT), and Natural Language Generation (NLG).
+Due to space limitations, we present the detailed results and analysis for NLG in Appendix A.3.
 The experimental setup covers the datasets, baseline models, and hyperparameter configurations to ensure reproducibility and comparability of the results.
@@
 We select several representative benchmarks in the current PEFT and small-model fine-tuning literature.
-For CSR, we evaluate on LLaMA-3.1-8B dubey2024llama using PIQA, BoolQ, HellaSwag, WinoGrande, SIQA, OBQA, and ARC-e/ARC-c, covering typical language reasoning scenarios such as physical commonsense, Boolean question answering, narrative completion, and scientific QA liu2024dora.
-For VIT, we evaluate LLaVA-1.5-7B liu2023visual; liu2024improved on ScienceQA, POPE, MMBench, GQA, VisWiz, VQAv2, and VQAT, which are generative language tasks with open-ended answers conditioned on images and textual instructions, rather than fixed-choice classification.
-For NLU, we build on DeBERTaV3-base hedebertav3 and use the GLUE benchmark datasets wang2018glue.
+For CSR, we evaluate on LLaMA-3.1-8B Dubey et al.
+(2024) using PIQA, BoolQ, HellaSwag, WinoGrande, SIQA, OBQA, and ARC-e/ARC-c, covering typical language reasoning scenarios such as physical commonsense, boolean question answering, narrative completion, and scientific QA Liu et al.
+(2024c).
+For VIT, we evaluate LLaVA-1.5-7B Liu et al.
+(2023, 2024a) on ScienceQA, POPE, MMBench, GQA, VisWiz, VQAv2, and VQAT, which are generative language tasks with open-ended answers conditioned on images and textual instructions.
+For NLU, we build on DeBERTaV3-base He et al.
+(2021) and use the GLUE benchmark datasets Wang et al.
+(2018).
 The GLUE tasks follow the official evaluation protocol: CoLA uses the Matthews correlation coefficient, STS-B uses correlation coefficients, and all other tasks use accuracy.
+Additionally, for the NLG tasks (detailed in Appendix), we evaluate Mistral-7B-v0.1 Jiang et al.
+(2023) on the Vicuna Chiang et al.
+(2023), LIMA Zhou et al.
+(2023) and Dolly Conover et al.
+(2023) datasets.
 Baselines.
-In our experiments, we select several mainstream parameter-efficient fine-tuning methods as baselines, covering two major categories: low-rank update methods and adapter-based methods.
-Specifically, LoRA, DoRA, and AdaLoRA are used to examine whether TS-PEFT can consistently provide additional gains under different low-rank update strategies, while Parallel Adapter hu2023llm is included to evaluate the method’s scalability on adapter-style architectures.
+In our experiments, we select several mainstream PEFT methods as baselines, covering two major categories: low-rank update methods and adapter-based methods.
+Specifically, LoRA, DoRA, and AdaLoRA are used to examine whether TS-PEFT can consistently provide additional gains under different low-rank update strategies, while Parallel Adapter Hu et al.
+(2023) is included to evaluate the method’s scalability on adapter-style architectures.
 For each baseline, we keep its original implementation, hyperparameters, and rank configuration unchanged, and apply TS-PEFT on top of it.
@@
 Training details.
-In all experiments, we freeze the backbone model and update only the PEFT parameters and the threshold $\tau$.
-We use the AdamW optimizer, where the PEFT parameters and $\tau$ share the same base learning rate, while the effective learning rate of $\tau$ is further modulated by the scaling factor s.
-For text tasks, we follow standard instruction-tuning configurations; for vision–language tasks, we adopt the original training pipeline of the corresponding models.
+In all experiments, we freeze the base model and update only the PEFT parameters and the threshold $\tau$.
+We use the AdamW optimizer, where PEFT is the default learning rate and the effective learning rate of $\tau$ is further modulated by the scaling factor $s$.
+For text tasks, we follow standard instruction-tuning configurations.
+For vision–language tasks, we adopt the original training pipeline of the corresponding models.
 Unless otherwise specified, we set $\alpha_{k}=1$, $\beta_{1}=0.9$, and $\beta_{2}=0.98$.
@@
 Under the same rank settings, we observe that TS-LoRA, TS-DoRA, and TS-AdaLoRA achieve performance comparable to or better than their respective baselines on most datasets.
-This indicates that selectively skipping updates on a subset of tokens does not weaken downstream performance; instead, it can help mitigate the interference caused by redundant updates.
-The table also reports the sparsity rate of TS-PEFT (i.e., the proportion of positions in the input sequence that are not modified by the PEFT output, as defined in Eq.(21)).
-Across different baselines and tasks, this proportion remains roughly within the 40%–60% range, meaning that fewer than half of the tokens actually receive PEFT updates.
+This indicates that selectively skipping updates on a subset of tokens does not weaken downstream performance.
+Instead, it can help alleviate the noise introduced by redundant updates.
+The table also reports the sparsity rate of TS-PEFT, which is the proportion of positions in the input sequence that are not modified by the PEFT output, as defined in Eq.(21).
+Across different baselines and tasks, this proportion remains roughly within the 50%–70% range, meaning that only 30% to 50% of the tokens actually receive PEFT updates.
 Even under this degree of sparsity, TS-PEFT is able to match or outperform the original baselines in commonsense reasoning, suggesting that standard PEFT exhibits substantial redundancy along the token dimension, and that threshold-based token-level selective updating provides a more efficient use of PEFT capacity.
@@
 Visual Instruction Tuning.
-For visual instruction tuning, we follow the training setup of liu2024dora, configuring LoRA and DoRA with rank 128 and setting the initial and target ranks of AdaLoRA to 256 and 128.
+For visual instruction tuning, we follow the training setup of Liu et al.
+(2024c), configuring LoRA and DoRA with rank 128 and setting the initial and target ranks of AdaLoRA to 256 and 128.
 Table 2 shows the results.
-Across all benchmarks, TS-LoRA, TS-DoRA, and TS-AdaLoRA yield consistent improvements of about 0.3–0.4 points, while operating under 50–65% token sparsity.
+In all benchmark tests, TS-LoRA, TS-DoRA, and TS-AdaLoRA achieve consistent improvements of about 0.3–0.4 points, while operating under 50%–70% token sparsity.
 This demonstrates that token-level redundancy is also present in multimodal generative tasks, and selective token updates remain effective, consistent with our observations in CSR.
@@
 The results show that TS-Adapter performs no worse than the original Adapter, suggesting that TS-PEFT is not bound to LoRA-style weight structures, but can function as a general token-level gating layer that can be stacked on diverse PEFT modules.
-Although TS-PEFT achieves a theoretical computation reduction of approximately 40%-60% in FLOPs within the adapter modules, current deep learning accelerators and libraries remain heavily optimized for dense matrix operations.
-Consequently, the imposed token-level sparsity currently serves as a theoretical computational bound and a probe for quantifying redundancy, rather than an immediate mechanism for wall-clock speedup.
-Nevertheless, with the emerging support for sparse tensor cores in next-generation hardware, TS-PEFT establishes a blueprint for translating this identified redundancy into tangible physical efficiency.
+TS-PEFT has achieved PEFT updates for only 30% - 70% of the tokens in multiple tasks, yet still maintain or even surpass the baseline performance.
+This clearly reveals the widespread token-level redundant updates in the current PEFT mode, and these updates also introduce noise that affects model performance.
+By quantifying this intrinsic redundancy, TS-PEFT establishes a clear blueprint for translating theoretical sparsity into substantial physical efficiency gains in the future.
 4.3 Analysis: Token Sparsity Reveals Model Adaptation Mechanism
-Figure 2: Performance of LLaMA3.1-8B on CSR benchmarks as the percentage of selected modules varies for each PEFT method.
-Note the performance drops (e.g., DoRA at 60%) when redundant, high-sparsity modules are forced to update, indicating noise injection.
 PIQA BoolQ HellaSwag WinoGrande SIQA OBQA ARC-e ARC-c Avg
@@
 Table 4: Performance of LLaMA3.1-8B on the commonsense reasoning task using various selection methods.
+Figure 2: Performance of LLaMA3.1-8B on CSR benchmarks as the percentage of selected modules varies for each PEFT method.
+Note the performance drops (e.g., DoRA at 60%) when redundant, high-sparsity modules are forced to update, indicating noise injection.
 Sparsity Patterns Reflect Architecture Bias.
-In LLaMA3.1-8B, there are 32 layers in total.
+There are 32 layers in LLaMA3.1-8B.
 We compute the sparsity of q_proj, k_proj, and v_proj in each layer and take their average to obtain the sparsity distribution of TS-PEFT across modules (detailed results are provided in Appendix A.5).
 Overall, the sparsity of TS-LoRA and TS-AdaLoRA is relatively evenly distributed across layers, with standard deviations of 9.8 and 11.4, respectively.
-In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, showing a more “polarized” pattern.
+In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, which shows a more polarized pattern.
 It should be noted that LoRA and AdaLoRA share the same weight structure despite having different rank allocation strategies, whereas DoRA adopts a distinct parameterization by decomposing weights into magnitude and direction.
@@
 Redundant Updates Introduce Optimization Noise.
-To further validate that the identified high-sparsity modules are indeed redundant and potentially harmful, we conducted a cumulative selection analysis.
-We expanded the set of target modules in ascending order of sparsity (i.e., adding modules from “critical” to “redundant”).
-As illustrated in Figure 2, the conventional wisdom that ”more trainable parameters imply better performance” does not hold.
+To further validate that the identified high-sparsity modules are indeed redundant and potentially harmful, we conduct a cumulative selection analysis.
+We expand the set of target modules in ascending order of sparsity.
+Specifically, we add the modules from the key ones to the redundant ones sequentially.
+As illustrated in Figure 2, the conventional wisdom that “more trainable parameters imply better performance” does not hold.
 For DoRA, performance peaks when only 50% of the modules are fine-tuned, while LoRA reaches its optimum at 80%.
-Remarkably, even utilizing just 20% of the modules yields results comparable to the optimal settings
-Crucially, we observe distinct “performance cliffs” when redundant modules are forced to update.
+Remarkably, even utilizing just 20% of the modules yields results comparable to the optimal settings.
+Crucially, we observe distinct performance cliffs when redundant modules are forced to update.
 For instance, DoRA suffers a sharp performance drop of 6.6% when the target proportion increases from 50% to 60%.
@@
 This phenomenon provides compelling evidence that updating modules characterized by high sparsity injects negative interference rather than meaningful adaptation.
-In contrast, AdaLoRA exhibits greater stability against these low-importance modules, likely due to its dynamic rank allocation mechanism which automatically limits the capacity assigned to redundant positions.
+In contrast, AdaLoRA exhibits greater stability against these low-importance modules, likely due to its dynamic rank allocation mechanism.
 5 Conclusion
-This paper revisits parameter-efficient fine-tuning from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers.
-We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 40–60% of token positions, yet consistently matches or surpasses strong baselines.
+This paper revisits PEFT from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers.
+We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 30%–70% of token positions, yet consistently matches or surpasses strong baselines.
 Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a theoretical basis for future hardware-software co-design targeting efficient sparse fine-tuning and inference.
@@
 All technical content, scientific claims, and experimental results were conceived, derived, and verified solely by the authors.
-A.2 The training loss progression with and without the Adaptive Moment Estimation
+A.2 Case Study
+Figure 3: Visualization of token-level sparsity learned by TS-LoRA on a BoolQ sample.
+To explicitly demonstrate the token selection of TS-PEFT, we visualize the learned token-level sparsity on a representative validation sample from the BoolQ dataset, as shown in Figure 3.
+By observing the distribution of the sparsity patterns, we find that TS-PEFT exhibits a clear semantic filtering mechanism when processing BoolQ tasks, which is manifested in three aspects:
+  • Suppression of High-Frequency Function Words and Syntactic Noise: The model automatically freezes high-frequency function words and stop words such as “is”, “that”, “the”, and “:”.
+These words already possess well-established syntactic functions from the pre-training phase, and forcibly fine-tuning them often introduces unnecessary optimization noise.
+TS-PEFT effectively preserves the linguistic knowledge of the pre-trained model by skipping these positions.
+  • Focus on Critical Domain Entities: For entity words closely related to the core logic of the question, such as “ethanol”, “energy”, and “produces”, the model chooses to update their parameters.
+This indicates that TS-PEFT can accurately isolate information carriers and concentrate the optimization budget thereon.
+  • Preservation of Generic Pre-trained Concepts: Interestingly, the model also tends to keep the core label words in the output format, such as “true”, “false”, frozen.
+This implies that the model can reuse the truth-value concepts established during pre-training without re-adapting them during fine-tuning.
+In conclusion, these qualitative results strongly support the hypotheses presented in the main text.
+We show that TS-PEFT is capable of acting as a dynamic probe that can autonomously identify and filter out redundant parts in the token level, thereby reducing the noise introduced by fine-tuning in inference tasks.
+A.3 Additional Results
+To further evaluate the generalization capability of TS-PEFT in open-ended text generation scenarios, we extend our experiments to NLG tasks.
+Specifically, we perform fine-tuning on the Alpaca dataset using the Mistral-7B-v0.1 model and subsequently evaluate performance on three representative benchmarks: Vicuna, LIMA, and Dolly.
+To measure generation quality comprehensively and objectively, we employed DeepSeek-R1-0325 as an automated evaluator, scoring model outputs on a 10-point scale across four dimensions: accuracy, relevance, fluency, and completeness.
+As shown in Table 5, variants of TS-PEFT consistently outperform their corresponding dense update baselines across all test sets.
+Specifically, TS-LoRA, TS-DoRA, and TS-AdaLoRA surpass their baselines by average scores of 0.11, 0.10, and 0.12, respectively, while maintaining a token-level sparsity of approximately 75%, that is, updating only about one-quarter of the tokens.
+This result not only demonstrates the robustness of TS-PEFT in generation tasks but also provides strong empirical support for the core hypothesis of this paper.
+That is, in the process of fine-tuning, updating a vast number of non-critical tokens is not only computationally redundant but may even introduce noise, whereas dynamically filtering and skipping these updates via TS-PEFT facilitates the generation of higher-quality and more logically coherent responses.
+Hyperparams Vicuna LIMA Dolly Avg
+LoRA $s=4\mathrm{e}{-5}$ $\lambda=1\mathrm{e}{-5}$ 6.78 7.13 8.00 7.30
+TS-LoRA 6.86 7.20 8.18 7.41
+Sparsity(%) 76.8 78.8 76.2 77.3
+DoRA $s=4\mathrm{e}{-5}$ $\lambda=6\mathrm{e}{-5}$ 7.02 6.95 7.82 7.26
+TS-DoRA 7.09 6.83 8.15 7.36
+Sparsity(%) 75.8 76.4 73.2 75.1
+AdaLoRA $s=4\mathrm{e}{-5}$ $\lambda=6\mathrm{e}{-5}$ 6.94 6.74 7.59 7.09
+TS-AdaLoRA 6.98 6.76 7.89 7.21
+Sparsity(%) 72.2 77.3 75.3 74.9
+Table 5: Evaluation scores of Mistral-7B-v0.1 on NLG benchmarks; the last row in each block shows the sparsity of TS-PEFT.
+A.4 The training loss progression with and without the Adaptive Moment Estimation
 Figure 4 illustrates the training dynamics of a single TS-PEFT layer (the v_proj in layers.20.self_attn of LLaMA3.1-8B) with and without Adam-style momentum on the threshold $\tau$.
-In both subfigures, the blue curve shows the training loss, while the red dashed curve tracks the corresponding threshold value over training steps (right-hand axis).
-When momentum is enabled (Figure 3(a)), the threshold increases smoothly and gradually saturates around a stable value, while the loss decreases rapidly at the beginning and then oscillates mildly around a low level.
+In both subfigures, the blue curve shows the training loss, while the red dashed curve tracks the corresponding threshold value over training steps.
+As shown in Figure 4 (a), when momentum is enabled, the threshold increases smoothly and gradually saturates around a stable value, while the loss decreases rapidly at the beginning and then oscillates mildly around a low level.
 This indicates that $\tau$ can adapt to the task in a stable manner without causing large fluctuations in the optimization trajectory.
 Figure 4: Training loss progression for layer base_model.layers.20.self_attn.v_proj of LLaMA3.1-8B.
-In contrast, when we remove the adaptive moment estimation and update $\tau$ using plain gradient descent (Figure 3(b)), the threshold trajectory becomes highly oscillatory and eventually drifts to a much larger range (above 0.6).
+In contrast, when we remove the adaptive moment estimation and update $\tau$ using plain gradient descent, as shown in Figure 4 (b), the threshold trajectory becomes highly oscillatory and eventually drifts to a much larger range (above 0.6).
 The loss curve also exhibits large, persistent oscillations and fails to converge to a similarly low level as in Figure 4(a).
-This unstable behavior at the layer level is consistent with the degradation observed on downstream CSR benchmarks such as BoolQ: the dev accuracy drops noticeably compared with the full TS-PEFT variant that uses Adam-style updates for $\tau$.
+This unstable behavior at the layer level is consistent with the degradation observed on downstream CSR benchmarks such as BoolQ.
+That is, the accuracy drops noticeably compared with the TS-PEFT that uses Adam-style updates for $\tau$.
 These results provide direct empirical evidence that adaptive momentum estimation is crucial for stabilizing the learning of token-level thresholds in practice.
-A.3 Module-wise average sparsity of LLaMA3.1-8B on the commonsense reasoning task
-Table 5 reports the detailed module-wise sparsity of TS-LoRA, TS-DoRA, and TS-AdaLoRA on LLaMA3.1-8B for commonsense reasoning.
-For each transformer layer (indexed from 0 to 31), we measure the token-level sparsity of the PEFT updates in the query, key, and value projections (q_proj, k_proj, and v_proj), and then compute their average (“avg” column).
+A.5 Module-wise average sparsity of LLaMA3.1-8B on the commonsense reasoning task
+Table 6 reports the detailed module-wise sparsity of TS-LoRA, TS-DoRA, and TS-AdaLoRA on LLaMA3.1-8B for CSR benchmarks.
+For each transformer layer, that is, indexed from 0 to 31, we measure the token-level sparsity of the PEFT updates in the query, key, and value projections (q_proj, k_proj, and v_proj), and then compute their average, the “avg” column in the figure.
 The last row summarizes the mean sparsity over all layers for each method.
-A dash (“–”) indicates that the corresponding module is not assigned any rank (i.e., effectively pruned by the underlying PEFT scheme).
-From the table, we observe that TS-LoRA and TS-AdaLoRA exhibit relatively smooth sparsity profiles across layers: their per-layer averages mostly stay within the 50%–70% range, and no single layer dominates the overall sparsity pattern.
+From the table, we observe that TS-LoRA and TS-AdaLoRA exhibit relatively smooth sparsity profiles across layers.
+Their per-layer averages mostly stay within the 50%–70% range, and no single layer dominates the overall sparsity pattern.
 This matches the moderate standard deviations reported in the main text (9.8 for TS-LoRA and 11.4 for TS-AdaLoRA), suggesting that these two methods tend to distribute token-level updates more evenly over depth.
 In contrast, TS-DoRA shows a much more polarized pattern: several layers (e.g., layers 0–2, 5–7, 11, 17, 24, 29) have near-100% sparsity in one or more projections, while a few layers (such as 8, 9, 14, 22) maintain comparatively low sparsity.
-This leads to a much larger standard deviation (41.3), indicating that TS-DoRA concentrates effective updates into a smaller subset of layers.
+This resulted in a relatively large standard deviation of 41.3, indicating that TS-DoRA concentrates effective updates into a smaller subset of layers.
 These fine-grained statistics support our observation in the main paper that the token-level sparsity learned by TS-DoRA can serve as a strong indicator of module importance and motivates the subsequent experiments that select only the low-sparsity (high-importance) modules for fine-tuning.
@@
 Table 6: Module-wise average sparsity (%) of LLaMA3.1-8B on commonsense reasoning. “-” denotes rank 0.
-A.4 Hyperparameters
-Across all experiments, we tune the sparsity-related hyperparameters s and $\lambda$ separately from the other training configurations.
-For CSR and VIT, we fix the scaling factor at s = $4\times 10^{-5}$ and select $\lambda$ from the interval $10^{-7},10^{-3}]$ based on development performance.
-For NLU tasks, we instead set $s=1\times 10^{-4}$ and choose $\lambda$ from $10^{-8},10^{-3}]$.
-The remaining hyperparameters, including rank, scaling coefficient, dropout rate, optimizer, learning rate, scheduler, batch size, number of epochs, and target modules, are summarized in Table 6.
+A.6 Hyperparameters
 Task Data Rank Scale Dropout Optimizer LR Scheduler Batch Epochs Target
@@
 NLU STS-B 4 0.5 0.2 AdamW 5e-4 Linear 32 45 Q,K,V,O,I
+NLG ALL 32 0.5 0.05 AdamW 2e-4 Linear 8 3 Q,K,V,O,Up,Down,Gate
 Table 7: Hyperparameter configuration for our experiments.
+Across all experiments, we tune the sparsity-related hyperparameters s and $\lambda$ separately from the other training configurations.
+For CSR and VIT, we fix the scaling factor at s = $4\times 10^{-5}$ and select $\lambda$ from the interval $10^{-7},10^{-3}]$ based on development performance.
+For NLU tasks, we instead set $s=1\times 10^{-4}$ and choose $\lambda$ from $10^{-8},10^{-3}]$.
+For NLG task, we set $s=4\times 10^{-5}$, and select $\lambda$ from $10^{-6},10^{-3}]$.
+The remaining hyperparameters, including rank, scaling coefficient, dropout rate, optimizer, learning rate, scheduler, batch size, number of epochs, and target modules, are summarized in Table 7.
+A.7 Hyperparameter Sensitivity Analysis
+Figure 5: Sensitivity analysis of the sparsity penalty term $\lambda$ for TS-DoRA on CSR benchmarks.
+The trend indicates that appropriate sparsity enhances performance by mitigating noise, whereas excessive sparsity impairs model capacity, thereby confirming the existence of an optimal level of redundancy.
+Impact of the Sparsity Penalty $\lambda$.
+To investigate the sensitivity of TS-PEFT to the sparsity regularization term, we conduct a study on the hyperparameter $\lambda$ using the TS-DoRA model on CSR benchmarks.
+As defined in the objective function Eq.
+(4), $\lambda$ governs the trade-off between minimizing the task loss $l(\tau)$ and enforcing sparsity.
+Figure 5 illustrates the trend of the Average Score on CSR tasks as $\lambda$ varies.
+We observe a distinct “inverted U-shaped” trend, which aligns well with our theoretical expectations:
+  • Low $\lambda$ Region (Left side of the curve): When $\lambda$ is extremely small, the penalty for activating tokens is negligible.
+Thus, the threshold $\tau$ remains at a low level, and the model behavior asymptotically converges to the standard, dense-update DoRA baseline.
+Although the performance is acceptable, it remains suboptimal due to the presence of redundant updates and optimization noise.
+  • Optimal $\lambda$ Region (Peak): As $\lambda$ increases to the optimal range, the model effectively prunes redundant token updates.
+This peak performance strongly validates our insight: selectively skipping non-critical updates can mitigate noise and enhance generalization capability.
+  • High $\lambda$ Region (Right side of the curve): When $\lambda$ is excessively large, the regularization term dominates the optimization objective.
+Even critical updates required for task adaptation are suppressed.
+This triggers underfitting and leads to a sharp decline in performance.
+Role of the Gradient Scaling Factor $s$.
+Regarding the hyperparameter $s$, we focus on its role within the threshold update rule defined in Eq.
+(20).
+Here, $s$ serves as a global scaling factor.
+Given that the derivative of the step function is theoretically zero almost everywhere, $s$ essentially determines the effective step size magnitude for learning $\tau$, functioning similarly to a Base Learning Rate.
+Empirically, we found that treating $s$ as a fixed constant suffices to achieve stable convergence.
+As shown in Appendix A.4, we set $s$ to $4\times 10^{-5}$ for CSR, VIT and NLG tasks and $1\times 10^{-4}$ for NLU tasks, relying on adaptive momentum estimation to manage dynamic adjustments during training.